### v1

# thanks: Corresponding authorthanks: Corresponding author

Abstract
Quantum steering is a crucial quantum resource that lies intermediate between entanglement and Bell nonlocality. Gaussian channels, meanwhile, play a foundational role in diverse quantum protocols, secure communication, and related fields. In this paper, we focus on several classes of Gaussian channels associated with quantum steering: Gaussian steering-annihilating channels, Gaussian steering-breaking channels, Gaussian unsteerable channels, and maximal Gaussian unsteerable channels. We give the concepts of these channels, derive the necessary and sufficient conditions for a Gaussian channel to belong to each class, and explore the intrinsic relationships among them. Additionally, since quantifying the steering capability of Gaussian channels in continuous-variable systems requires an understanding of the structure of free superchannels, we also provide a detailed characterization of Gaussian unsteerable superchannels and maximal Gaussian unsteerable superchannels.
preprint: APS/123-QED
I Introduction
[…]
Gaussian states are a special class of quantum states in continuous-variable (CV) systems, playing a pivotal role in quantum optics and quantum information theory (SLP ; WPG1 ; S1 ). Over the past few years, the EPR steering criteria and measures for Gaussian states have garnered considerable attention from researchers (see, e.g., JiL ; KSA ; KLRA ; ZYL ; TGRF ; MKIB ; BJC ; YHQ and the references therein). Notably, quantum information processing inevitably involves quantum channels. As a distinctive class of quantum channels, Gaussian quantum channels not only furnish a core theoretical framework for elucidating the inherent physical limitations of quantum communication and quantum computing, but also directly underpin the translation of quantum technologies from theoretical concepts to practical applications (BSS ; MDG ; CEGH ; CD ; NGA ; LD ; PMGH ). They thus hold irreplaceable significance in quantum systems, especially in optical quantum systems. One primary objective of this paper is to investigate three types of special Gaussian quantum channels with respect to EPR steering: Gaussian steering-annihilating channels that completely eliminate steering; Gaussian steering-breaking channels that locally disrupt steering; and maximal Gaussian unsteerable channels that map Gaussian unsteerable states into Gaussian unsteerable states.
Regarding the correlation measures and resource theories of quantum channels, substantial research efforts have been devoted. Bäuml et al. BDW proposed several entanglement measures tailored for bipartite quantum channels. Mani MK introduced the concepts of cohering and decohering power of quantum channels, along with corresponding quantification methods. Xu X19 established a coherence resource theory for channels in finite-dimensional systems, while the authors of YX1 developed a general operational resource theory framework for quantum channels in such systems. For Gaussian channel resource theories, Xu X11 constructed a coherence resource theory specific to Gaussian channels and proposed a coherence measure for them, based on the relative entropy coherence measure for Gaussian states. Recall that a quantum channel resource theory is defined as a tuple $(\mathcal{F},\mathcal{O},\mathcal{R})$, where $\mathcal{F}$ is the set of free channels that do not have any resource, $\mathcal{O}$ is the set of free superoperations which transform free channels into free channels; and $\mathcal{R}$ is the set of channel resource measures which map quantum channels into nonnegative real numbers satisfying the following two fundamental conditions:
$(f_{1})$ non-negativity: $\mathcal{R}\left(\phi\right)\geq 0$ for all $\phi\in\mathcal{C}(H)$ (the set of all quantum channels on a separable complex Hilbert space $H$), and $\mathcal{R}\left(\phi\right)=0$ for any $\phi\in{\mathcal{F}}$;
$(f_{2})$ monotonicity: $\mathcal{R}\left({\Psi\left(\phi\right)}\right)\leq\mathcal{R}\left(\phi\right)$ holds for all $\phi\in{\mathcal{F}}$ and all $\Psi\in{\mathcal{O}}$.
To lay the foundation for the future development of a steering resource theory for Gaussian channels, another core objective of this paper is to investigate the structure of free superchannels, specifically Gaussian unsteerable superchannels and maximal Gaussian unsteerable superchannels.
This paper is structured as follows. In Section II, we review fundamental concepts related to continuous-variable (CV) systems, including Gaussian states, Gaussian channels, Gaussian unsteerable channels, and Gaussian quantum steering. In Section III, we formally define Gaussian steering-annihilating channels and Gaussian steering-breaking channels, analyze their structural properties, derive the necessary and sufficient conditions for a Gaussian channel to be classified as either type, and explore the relationships between these channels. Section IV is dedicated to characterizing Gaussian unsteerable superchannels and maximal Gaussian unsteerable superchannels. Section V presents a concise summary of the work.
II Preliminaries
In this section, we briefly recall some notions and notations about Gaussian states and Gaussian quantum channels.
II.1 Gaussian states
Consider an $N$-mode CV system with state space $H=H_{1}\otimes H_{2}\otimes\cdots\otimes H_{N}$, where each $H_{k}\ (1\leq k\leq N)$ is an infinite-dimensional separable complex Hilbert space. Denote by $\mathcal{S}(H)$ the set of all quantum states (that is, positive bounded linear operators with trace 1) on $H$.
For any state $\rho\in\mathcal{S}(H)$, its characteristic function $\chi_{\rho}$ is defined as
$\chi_{\rho}(z)={\rm tr}(\rho W(z)),$
where $z=(x_{1},y_{1},\cdots,x_{N},y_{N})^{\rm T}\in{\mathbb{R}}^{2N}$, $W(z)=\exp(i{R}^{\rm T}z)$ is the Weyl displacement operator, ${R}=(R_{1},R_{2},\cdots,R_{2N})=(\hat{Q}_{1},\hat{P}_{1},\cdots,\hat{Q}_{N},\hat{P}_{N})$, $\hat{Q}_{k}=(\hat{a}_{k}+{\hat{a}_{k}}^{\dagger})/\sqrt{2}$ and $\hat{P_{k}}=-i({\hat{a}_{k}}-{\hat{a}_{k}}^{\dagger})/\sqrt{2}$ ($k=1,2,\cdots,N$) are respectively the position and momentum operators. Here, $\hat{a}_{k}^{\dagger}$ and $\hat{a}_{k}$ are the creation and annihilation operators in the $k$th mode satisfying the Canonical Commutation Relation (CCR):
$\hat{a}_{k},\hat{a}_{l}^{\dagger}]=\delta_{kl}I\ {\rm and}\ [\hat{a}_{k}^{\dagger},\hat{a}_{l}^{\dagger}]=[\hat{a}_{k},\hat{a}_{l}]=0,\ k,l=1,2,\cdots,N.$
Particularly, $\rho$ is called a Gaussian state if $\chi_{\rho}(z)$ is of the form
$\chi_{\rho}(z)=\exp[-\frac{1}{4}z^{\rm T}\Gamma z+i{\mathbf{d}}^{\rm T}z],$
where
$\begin{array}[]{rl}{\mathbf{d}}=&(\langle\hat{R}_{1}\rangle,\langle\hat{R}_{2}\rangle,\ldots,\langle\hat{R}_{2N}\rangle)^{\rm T}\\ =&({\rm tr}(\rho R_{1}),{\rm tr}(\rho R_{2}),\ldots,{\rm tr}(\rho R_{2N}))^{\rm T}\in{\mathbb{R}}^{2N}\end{array}$
is called the mean or the displacement vector of $\rho$ and $\Gamma=(\gamma_{kl})\in\mathcal{M}_{2N}(\mathbb{R})$ is called the covariance matrix (CM) of $\rho$ defined by $\gamma_{kl}={\rm tr}[\rho(\Delta\hat{R}_{k}\Delta\hat{R}_{l}+\Delta\hat{R}_{l}\Delta\hat{R}_{k})]$ with $\Delta\hat{R}_{k}=\hat{R}_{k}-\langle\hat{R}_{k}\rangle$ (SLP ). Here, $\mathcal{M}_{d}(\mathbb{R})$ stands for the algebra of all $d\times d$ matrices over the real field $\mathbb{R}$. So, any Gaussian state $\rho$ with CM $\Gamma$ and displacement vector $\mathbf{d}$} will sometimes be represented as $\rho(\Gamma,{\mathbf{d}})$. Note that $\Gamma$ is real symmetric and satisfies the condition
$\Gamma+i\Omega_{N}\geq 0,\ \ {\rm where}\ \ \Omega_{N}=\oplus_{k=1}^{N}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
Now, divide the $N$-mode CV system into $m$-mode CV subsystem A and $n$-mode CV subsystem B, with state space $H=H_{A}\otimes H_{B}$ and $N=m+n$. Assume that $\rho$ is any $(m+n)$-mode bipartite Gaussian state. Then its CM $\Gamma_{\rho}$ can be written as
$\Gamma_{\rho}=\left(\begin{array}[]{cc}A&C\\ C^{\rm T}&B\end{array}\right),$ (3)
where $A\in\mathcal{M}_{2m}({\mathbb{R}}),\ B\in\mathcal{M}_{2n}({\mathbb{R}}),C\in\mathcal{M}_{2m\times 2n}({\mathbb{R}})$. Particularly, if $n=m=1$, then $\Gamma$ has the following standard form:
[…]
where $a,b\geq 1,\ ab-c^{2}\geq 1,\ ab-d^{2}\geq 1$. For more details about Gaussian states, see WPG1 ; S1 .
In the rest of this paper, if there is no explanation, we always assume that $H$, $H_{A}$ and $H_{B}$ are infinite-dimensional separable complex Hilbert spaces with $H=H_{A}\otimes H_{B}$.
II.2 EPR steering
A measurement assemblage ${\mathcal{MA}}=\{M_{a|x}\}_{a,x}$ is a collection of positive operators $M_{a|x}\geq 0$ satisfying $\sum_{a}M_{a|x}=I$ for each $x$. Such a collection represents one positive-operator-valued measurement (POVM), describing a general quantum measurement, for each $x$. In a (bipartite) EPR steering scenario, one party performs measurements on a shared state $\rho_{AB}\in{\mathcal{S}}(H_{A}\otimes H_{B})$, which steers the quantum state of the other particle. If Alice performs a set of measurements $\{M^{A}_{a|x}\}_{a,x}$, then the collection of sub-normalized of Bob is an assemblage $\{\rho^{B}_{a|x}\}_{a,x}$ with
$\rho^{B}_{a|x}={\rm Tr}_{A}((M^{A}_{a|x}\otimes I_{B}){\rho}_{AB}).$
If every assemblage on Bob $\{\rho^{B}_{a|x}\}_{a,x}$ can be explained by a local hidden state (LHS) model:
$\rho^{B}_{a|x}=\sum_{\lambda}p_{\lambda}p(a|x,\lambda)\sigma_{\lambda},$
where $\lambda$ is a hidden variable, distributed according to $p_{\lambda}$, $\sigma_{\lambda}$ are “hidden states” of Bob, and $p(a|x,\lambda)$ are local “response functions” of Alice, then we say that $\rho_{AB}$ has a LHS form, or does not demonstrate steering (WJ ). Otherwise, if there exist measurements such that $\rho^{B}_{a|x}$ does not admit such LHS decomposition, then $\rho_{AB}$ is called steerable from A to B. Symmetrically, we can define the steerability of $\rho_{AB}$ from B to A. Here, we point out that, if there is no special illustration, we focus on the steerability from A to B in this paper.
In CV systems, Gaussian POVM (GPOVM) plays an important role. Recall that an $N$-mode GPOVM $\Pi=\{\Pi(\alpha)\}$ is defined as
$\Pi(\alpha)=\frac{1}{\pi^{N}}D(\alpha)\varpi D^{\dagger}(\alpha),$
where $\alpha=(\alpha_{1},\cdots,\alpha_{N})^{\rm T}\in{\mathbb{C}^{N}}$, $D(\alpha)={\rm exp}[\sum_{j=1}^{N}(\alpha_{j}\hat{a}_{j}^{\dagger}-\alpha_{j}^{*}\hat{a}_{j})]$ is the $N$-mode Weyl displacement operator and $\varpi$ is a zero mean $N$-mode Gaussian state, which is called the seed state of $\Pi$ (GC ).
For any bipartite Gaussian state, the authors in WJ derived a linear matrix inequality steering criterion via GPOVMs. Assume that $\rho\in{\mathcal{S}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3). As demonstrated in WJ , $\rho$ is unsteerable by the subsystem A’s all GPOVMs if and only if
$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0,\ \ {\rm where}\ \ \Omega_{n}=\oplus_{k=1}^{n}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$. (5)
Obviously,
$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\Leftrightarrow\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n}).$
Denote respectively by ${\mathcal{GS}}(H)$ and ${\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ the set of all Gaussian states on $H$ and the set of all Gaussian unsteerable states from A to B on $H_{A}\otimes H_{B}$, that is,
$\mathcal{GS}(H)=\{\rho\in{\mathcal{S}}(H):\ \rho\ \mbox{is a Gaussian state}\}$
and
$\begin{array}[]{rl}&{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =&\{\rho\in{\mathcal{GS}}(H_{A}\otimes H_{B}):\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\}.\end{array}$
II.3 Gaussian unsteerable channels
Recall that a Gaussian channel is a quantum channel which transforms Gaussian states into Gaussian states (PMGH ; CEGH ). An $N$-mode Gaussian channel $\phi$ on $H$ can be described by $\phi=\phi\left({K,M,\mathbf{d}}\right)$, which acts on $\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in\mathcal{GS}(H)$ as
$\mathbf{d}_{\rho}\mapsto K\mathbf{d}_{\rho}+\mathbf{d},\ \ \Gamma_{\rho}\mapsto K\Gamma_{\rho}K^{\rm T}+M,$ (6)
$\mathbf{d}\in\mathbb{R}^{2N}$ is a column displacement vector, $K,M\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $M=M^{\rm T}$ and the completely positive condition
$M+i\Omega_{N}-iK\Omega_{N}{K^{\rm T}}\geq 0.$ (7)
Denote by ${\mathcal{GC}}(H)$ the set of all Gaussian channels on $H$. Here, we give three known Gaussian channels which are used frequently.
Attenuator channels. A single-mode attenuator channel $\phi_{\theta}^{n_{th}}(K,M,\mathbf{d})$ is a deterministic Gaussian channel with
[…]
where $\theta\in[0,2\pi]$ and thermal noise $n_{th}\geq 1$. Particularly, if $n_{th}=1$, then the channel is called a pure lossy channel.
Constant channels. A Gaussian channel $\Theta(K,M,\mathbf{d})\in{\mathcal{GC}}(H)$ is called a constant channel if there exists some $\rho_{0}(\Gamma_{0},{\mathbf{d}}_{0})\in\mathcal{GS}(H)$ such that $\Theta(\rho)=\rho_{0}$ for all $\rho\in\mathcal{GS}(H)$. In this case, $\Theta(K,M,\mathbf{d})$ can be represented as $\Theta(K,M,\mathbf{d})=\Theta(0,\Gamma_{0},\mathbf{\mathbf{d}}_{0})$.
Identity Gaussian channel. A Gaussian channel $\psi(K,M,\mathbf{d})\in{\mathcal{GC}}(H)$ is called an identity channel if $\psi(\rho)=\rho$ holds for all $\rho\in\mathcal{GS}(H)$. In this case, $K=I$, $M=0$ and $\mathbf{d}=0$.
In YHQ , the authors gave the definition of Gaussian unsteerable channels. Any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}(H_{A}\otimes H_{B})$ is called Gaussian unsteerable (from A to B) if $K$ and $M$ satisfy the following relation
$M+({0_{2m}}\oplus i{\Omega_{n}})-K\left({0_{2m}}\oplus i{\Omega_{n}}\right){K^{\rm T}}\geq 0;$ (9)
and is called maximal Gaussian unsteerable (from A to B) if $\phi({\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})$. Let
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =&\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}(H_{A}\otimes H_{B}):\\ &M+(0_{2m}\oplus i{\Omega_{n}})-K\left(0_{2m}\oplus i{\Omega_{n}}\right){K^{\rm T}}\geq 0\}\end{array}$
and
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\\ =&\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}(H_{A}\otimes H_{B}):\\ &\phi({\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})\}.\end{array}$
It is shown in YHQ that the set ${\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ is a proper (but) large subset of ${\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$. Obviously, the $(m+n)$-mode identity Gaussian channel $\phi(I,0,0)\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
II.4 Gaussian superchannels
Recall that a superchannel is a completely positive linear map transforming any quantum channels into quantum channels; and a Gaussian superchannel is a superchannel transforming any Gaussian channels into Gaussian channels (X11 ). Write
${\mathcal{SGC}}(H)=\{\mbox{all Gaussian superchannels on }H\}.$
Xu X11 gave two different representations of Gaussian superchannels.
Theorem 1.
(X11 ) Any $N$-mode Gaussian superchannel $\Phi\in{\mathcal{SGC}}(H)$ can be represented by $\Phi\left({A,E,Y,\nu}\right)$, where $\nu\in{\mathbb{R}}^{2N}$, $A,E,Y\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $Y=Y^{\rm T}$, $EE^{\rm T}={I_{2N}}$ and
$Y+i\Omega_{N}-iA\Omega_{N}A^{\rm T}\geq 0,\ \ i\Omega_{N}-iE\Omega_{N}E^{\rm T}\geq 0.$ (10)
Moreover, for any $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}(H)$, we have
(1) $\Phi\left({\phi\left({K,M,\mathbf{d}}\right)}\right)=\phi^{\prime}\left({K^{\prime},M^{\prime},\mathbf{d}^{\prime}}\right)$ with
$K^{\prime}=AK\Sigma_{N}E^{\rm T}\Sigma_{N},\ M^{\prime}=AMA^{\rm T}+Y,\ \mathbf{d}^{\prime}=A\mathbf{d}+\nu,$
[…]
(2) $\Phi\left(\phi\left({K,M,\mathbf{d}}\right)\right)={\phi_{2}}\circ\phi\circ{\phi_{1}}$ with some ${\phi_{1}}\left(K_{1},M_{1},\mathbf{d}_{1}\right)$ and ${\phi_{2}}\left(K_{2},M_{2},\mathbf{d}_{2}\right)\in{\mathcal{GC}}(H)$, where $K_{1}=\Sigma_{N}E^{\rm T}\Sigma_{N},$ $M_{1}=0,$ ${\mathbf{d}_{1}}=0,$ $K_{2}=A,$ $M_{2}=Y$, ${\mathbf{d}_{2}}=\upsilon.$
III Gaussian steering-annihilating and Gaussian steering-breaking channels
In this section, we will first give two concepts of Gaussian steering-annihilating channels and Gaussian steering-breaking channels, and then discuss their properties.
Definition 2.
Assume that $H,K$ are any infinite-dimensional separable complex Hilbert spaces. We say that any Gaussian channel $\phi\in{\mathcal{GC}}(H\otimes K)$ is Gaussian steering-annihilating if it sends all Gaussian states into Gaussian unsteerable states, that is, if $\phi(\mathcal{GS}(H\otimes K))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H\otimes K)$; and any Gaussian channel $\psi\in\mathcal{GC}(H)$ is Gaussian steering-breaking if $(\psi\otimes I_{K})(\rho)$ is always unsteerable for all Gaussian states $\rho\in{\mathcal{GS}}(H\otimes K)$, that is, if $(\psi\otimes I_{K})(\mathcal{GS}(H\otimes K))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H\otimes K)$.
Let
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\\ =\{&\mbox{all Gaussian steering-annihilating }\\ &\mbox{channels on }H_{A}\otimes H_{B}\}\end{array}$
and
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{SB}}(H)\\ =&\{\mbox{ all Gaussian steering-breaking channels on }H\}.\end{array}$
By Definition 2, the following useful property is obvious.
Proposition 3.
If $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$, and $\psi\in{\mathcal{GC}}(H_{A}\otimes H_{B})$, then $\phi\circ\psi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$; if $\phi\in{\mathcal{GC}}_{\mathcal{SB}}(H)$ and $\psi\in{\mathcal{GC}}(H)$, then both $\phi\circ\psi$ and $\psi\circ\phi$ belong to ${\mathcal{GC}}_{\mathcal{SB}}(H)$.
Next, we first give a sufficient condition for Gaussian channels being Gaussian steering-annihilating.
Theorem 4.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any $(m+n)$-mode channel. If $\phi$ satisfies the condition
$M+({0_{2m}}\oplus i{\Omega_{n}})-iK\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\geq 0,$ (11)
then $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$.
Proof.
Assume that $\rho\in{\mathcal{GS}}\left(H_{A}\otimes H_{B}\right)$ is any Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3). Then $\phi(\rho)$ has the CM $\Gamma_{\phi(\rho)}=K\Gamma_{\rho}{K^{\rm T}}+M$. As $\Gamma_{\rho}+i\left({{\Omega_{m}}\oplus{\Omega_{n}}}\right)\geq 0$, by the assumption (11), we have
$\Gamma_{\phi(\rho)}+(0_{2m}\oplus i{\Omega_{n}})$
$=$ $K\Gamma_{\rho}{K^{\rm T}}+M+(0_{2m}\oplus i{\Omega_{n}})$
$\geq$ $K(-i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T}+M+(0_{2m}\oplus i\Omega_{n})\geq 0.$
It follows that $\phi\left(\rho\right)\in{\mathcal{GS}}_{\mathcal{U}S}(H_{A}\otimes H_{B})$. So $\phi$ is Gaussian steering-annihilating. ∎
Notice that the condition (11) is only sufficient but not necessary for a channel being steering-annihilating. In fact, there exist Gaussian steering-annihilating channels which do not satisfy the condition (11).
Example 5.
Take a $(1+1)$-mode Gaussian channel $\phi_{1}=\phi(K_{1},M_{1},\mathbf{d_{1}})$, where $M_{1}=I_{4}$ and
[…]
Then $\phi_{1}$ is steering-annihilating, but does not satisfy the condition (11).
In fact, it is easily checked that $\phi\left({{K_{1}},{M_{1}},{\rm d_{1}}}\right)$ satisfies the condition
[…]
but
[…]
So $\phi$ does not satisfy the condition (11).
However, for any $(1+1)$-mode Gaussian state $\rho$ with the standard CM $\Gamma_{\rho}$ in Eq.(4), by a numerical calculation, one can obtain
[…]
which implies that $\phi$ is steering-annihilating.
While Ineq.(11) is not a necessary condition, the subsequent Ineq.(12) provides a sufficient and necessary condition for a Gaussian channel becoming steering-annihilating.
Theorem 6.
Any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$ if and only if
$\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$ (12)
$\geq$ $|\mathbf{w}^{{\dagger}}({0_{2m}}\oplus{\Omega_{n}})\mathbf{w}|$
holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
Proof.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any Gaussian channel. If $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$, then for any $\rho\in{\mathcal{GS}}\left(H_{A}\otimes H_{B}\right)$ with CM $\Gamma_{\rho}$, by Eqs.(5)-(6), one has
$K\Gamma_{\rho}K^{\rm T}+M\geq\pm i(0_{2m}\oplus\Omega_{n}),$
and so
$\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}$
holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
Note that $\Gamma_{\rho}\geq\pm i(\Omega_{m}\oplus\Omega_{n})$, and $\inf\limits_{\sigma\geq\pm i\Omega_{N}}{\bf w}^{\dagger}\sigma{\bf w}=|{\bf w}^{\dagger}\Omega_{N}{\bf w}|$ for all $N$-mode quantum states $\sigma$ PMGH . Thus, the above inequality implies
$|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
for all $\mathbf{w}\in\mathbb{C}^{2(m+n)},$ that is, Ineq.(12) holds.
On the other hand, if Ineq.(12) holds, then for any $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, one gets
$\mathbf{w}^{{\dagger}}\Gamma_{\phi(\rho)}\mathbf{w}$ $=$ $\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
$\geq$ $\inf_{\Gamma_{\rho}\geq\pm i\left({\Omega_{m}\oplus{\Omega_{n}}}\right)}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
$=$ $|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}$
$\geq$ $|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
$\geq$ $\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.$
So $\Gamma_{\phi(\rho)}\geq\pm i(0_{2m}\oplus\Omega_{n})$, which means that $\phi(\rho)$ is unsteerable. It follows that $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$. ∎
Note that
$\left|{{\varphi_{r}}}\right\rangle=\frac{1}{\cosh r}{\sum\limits_{j=0}^{\infty}{\left({\tanh r}\right)}^{j}}\left|j_{H}\right\rangle\left|j_{H}\right\rangle\in H\otimes H$
is a Gaussian squeezed pure state, where $\{|j_{H}\rangle\}$ is the Fock basis of $H$ and $r\in\mathbb{R}$ is the squeezed parameter. For any Gaussian channel $\phi\in{\mathcal{GC}}(H)$, define
${\rho_{\phi}}=(\phi\otimes I_{H})(\left|{{\varphi_{r}}}\right\rangle\left\langle{{\varphi_{r}}}\right|).$
Clearly, $\rho_{\phi}\in{\mathcal{GS}}(H\otimes H)$. By the unsteerability of $\rho_{\phi}$, we can give necessary and sufficient conditions for Gaussian steering-breaking channels.
Theorem 7.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H\right)$ is any N-mode Gaussian channel. Then the following statements are equivalent.
(1) $\phi\in{\mathcal{GC}}_{\mathcal{SB}}(H)$ is Gaussian steering-breaking.
(2) ${\rho_{\phi}}\in{\mathcal{GS}}_{\mathcal{US}}(H\otimes H)$.
(3) The matrices $K$ and $M$ satisfy the condition
$M-iK\Omega_{N}K^{\rm T}\geq 0.$
Proof.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H\right)$ is any $N$-mode Gaussian channel.
$\left(1\right)\Rightarrow\left(2\right)$: By Definition 2, this is obvious.
$\left(2\right)\Rightarrow\left(3\right)$: Assume that ${\rho_{\phi}}\in{\mathcal{GS}}_{\mathcal{US}}(H\otimes H)$. Note that the CM $\Gamma_{{\rho_{\phi}}}$ of $\rho_{\phi}$} has the form X11
[…]
and
$\Gamma_{{\rho_{\phi}}}+(0_{2N}\oplus i{\Omega_{N}})\geq 0.$ (14)
It is well known that a Hermitian matrix $W=\left(\begin{array}[]{cc}W_{11}&W_{12}\\ W_{12}^{{\dagger}}&W_{22}\end{array}\right)\geq 0$ if and only if $W_{22}\geq 0$ and $W_{11}-W_{12}W_{22}^{-1}W_{12}^{{\dagger}}\geq 0$ (ZH ). Then Ineq.(14) implies
$\begin{cases}\cosh 2rI_{2N}+i{\Omega_{N}}\geq 0,\\ \begin{array}[]{rl}&\cosh 2rK{K^{\rm T}}+M\\ &-\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N}+i\Omega_{N})^{-1}\Sigma_{N}{K^{\rm T}}\geq 0.\end{array}\end{cases}$
Note that $(T+cI)^{-1}=\frac{1}{c}(I-\frac{T}{c})+o(c^{-2})$ (S1 ), where $o(c^{-2})$ stands for the Landau little which will be neglected when taking to $c\to\infty$. So
$0\leq\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N}+i{\Omega_{N}})^{-1}\Sigma_{N}K^{\rm T}]$
$=$ $\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-\sinh^{2}(2r)K\Sigma_{N}[\frac{1}{\cosh 2r}(I_{2N}-\frac{i\Omega_{N}}{\cosh 2r})+o(\frac{1}{\cosh 2r})]\Sigma_{N}{K^{\rm T}}]$
$=$ $\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-K\Sigma_{N}[\frac{\sinh^{2}(2r)}{\cosh 2r}I_{2N}-i\tanh^{2}(2r){\Omega_{N}}+\sinh^{2}(2r)o(\frac{1}{\cosh 2r})]\Sigma_{N}K^{\rm T}]=M-iK{\Omega_{N}}{K^{\rm T}}.$
[…]
Since $\Gamma_{\rho}+i({\Omega_{N}}\oplus{\Omega_{N^{\prime}}})\geq 0$, a direct calculation gives
$Y+i{\Omega_{N^{\prime}}}\geq 0$}
and
$X+i{\Omega_{N}}-Z\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}{Z^{\rm T}}\geq 0$}.
This implies
$KX{K^{\rm T}}+M-KZ{\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}}{Z^{\rm T}}{K^{\rm T}}$
$=$ $K\left({X-Z{{\left({Y+i{\Omega_{N^{\prime}}}}\right)}^{-1}}{Z^{\rm T}}}\right){K^{\rm T}}+M$
$\geq$ $K\left({-i{\Omega_{N}}}\right){K^{\rm T}}+M=M-iK{\Omega_{N}}{K^{\rm T}}\geq 0.$
Also note that
$KX{K^{\rm T}}+M\geq 0$
by the denifition of the channel $\phi$. It follows that ${\Gamma_{\left({\phi\otimes I_{H^{\prime}}}\right)(\rho)}}+i(0_{N}\oplus{\Omega_{N^{\prime}}})\geq 0$. So $(\phi\otimes{I_{H^{\prime}}})\left(\rho\right)$ is unsteerable, and hence $\phi$ is steering-breaking. ∎
In the end of this section, we will investigate the relationship between Gaussian steering-breaking channels, Gaussian steering-annihilating channels and maximal Gaussian unsteerable channels.
By their definitions, it is obvious that
${\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\subset{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}).$
As a consequence of Proposition 3, for any $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$ and $\psi\in{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})$, we have
$\phi\circ\psi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\cap{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B}),$
which means that there are Gaussian channels which are simultaneously steering-breaking and steering-annihilating.
Next, take an $(m+n)$-mode constant channel $\Theta=\Theta(0,\Gamma_{0},\mathbf{d}_{0})$. Obviously, by Theorem 7, $\Theta$ is steering-breaking. However, if $\Gamma_{0}+(0_{2m}\oplus i\Omega_{n})\ngeq 0$, $\Theta$ is neither steering-annihilating nor maximal Gaussian unsteerable. So
${\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$
and
${\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}).$
There also exist Gaussian steering-annihilating channels which are not Gaussian steering-breaking, that is,
${\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B}).$
For example, take $\tilde{\phi}=\phi\otimes I_{B}\in{\mathcal{GC}}(H_{A}\otimes H_{B})$, where $\phi=\phi(\cos\theta I_{2},\sin^{2}\theta I_{2},\mathbf{d}_{0})$ is a single-mode attenuator channel. Then $\tilde{\phi}=\tilde{\phi}(K,M,\mathbf{d})$ can be represented by
[…]
It can be easily checked that, when $0\leq\cos\theta\leq\frac{\sqrt{2}}{2}$,
$M+({0_{2}}\oplus i{\Omega_{1}})-iK\left(\Omega_{1}\oplus\Omega_{1}\right){K^{T}}$
$=$ $\left(\begin{array}[]{cccc}\sin^{2}\theta&-i\cos^{2}\theta&0&0\\ i\cos^{2}\theta&\sin^{2}\theta&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)\geq 0,$
but
$M-iK{\Omega_{2}}K^{T}=M-iK\left(\Omega_{1}\oplus\Omega_{1}\right)K^{T}$
$=$ $\left(\begin{array}[]{cccc}\sin^{2}\theta&-i\cos^{2}\theta&0&0\\ i\cos^{2}\theta&\sin^{2}\theta&0&0\\ 0&0&0&-i\\ 0&0&i&0\end{array}\right)\ngeq 0.$
Thus, by Theorem 4 and Theorem 7, $\tilde{\phi}$ is steering-annihilating, but not steering-breaking.
For the relation of these Gaussian channels, see Fig.1.
Fig. 1. The relationship between Gaussian steering-breaking channels, Gaussian steering-annihilating channels and maximal Gaussian unsteerable channels.
IV Gaussian unsteerable superchannels
In this section, we discuss two special type of Gaussian unsteerable superchannels, that is, maximal Gaussian unsteerable superchannels and Gaussian unsteerable superchannels. Here, we say that a Gaussian superchannel is Gaussian unsteerable if it maps any Gaussian unsteerable channels into Gaussian unsteerable channels; and is maximal Gaussian unsteerable if it maps any maximal Gaussian unsteerable channels into maximal Gaussian unsteerable channels.
Denote by ${\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ the set of all maximal Gaussian unsteerable superchannels, that is,
$\begin{array}[]{rl}&{\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\\ =\{&\Phi\in{\mathcal{SGC}}(H_{A}\otimes H_{B}):\\ &\Phi({\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\},\end{array}$
and by ${\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ the set of all Gaussian unsteerable superchannels, that is,
$\begin{array}[]{rl}&{\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =\{&\Phi\in{\mathcal{SGC}}(H_{A}\otimes H_{B}):\\ &\Phi({\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\}.\end{array}$
To characterize maximal Gaussian unsteerable superchannels, it is necessary to discuss the structure of maximal Gaussian unsteerable channels. By a similar argument to that of Theorem 6, we can obtain a necessary and sufficient condition for channels to be maximal unsteerable.
Theorem 8.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any $(m+n)$-mode Gaussian channel. Then $\phi\in{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ if and only if
$\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(0_{2m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$
$\geq$ $|\mathbf{w}^{{\dagger}}({0_{2m}}\oplus{\Omega_{n}})\mathbf{w}|$
holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
Suppose that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian superchannel. By Theorem 1, there exist some ${\chi_{1}}\left(\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},0,0\right),{\chi_{2}}\left(A,Y,\nu\right)\in{\mathcal{GC}}(H_{A}\otimes H_{B})$ such that $\Phi\left(\phi\right)={\chi_{2}}\circ\phi\circ\chi_{1}$ for all $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}(H_{A}\otimes H_{B})$; and moreover, $\Phi\left(\phi\left(K,M,\mathbf{d}\right)\right)=\phi^{\prime}\left(K^{\prime},M^{\prime},\mathbf{d}^{\prime}\right)$ with
$K^{\prime}=AK\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},\ \ M^{\prime}=AMA^{\rm T}+Y.$
Take any $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ and any $\rho=\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in{\mathcal{GS}_{\mathcal{US}}(H_{A}\otimes H_{B})}$. Notice that the CM $\Gamma_{\tau}$ of any $(m+n)$-mode unsteerable Gaussian state $\tau$ can be written as $\Gamma_{\tau}=0_{2m}\oplus Q_{\tau}+P_{\tau}$ with some real matrix $Q_{\tau}\geq i{\Omega_{n}}$ and $P_{\tau}\geq 0$ MKIB . Now, if both $\chi_{1}$ and $\chi_{2}$ are maximal unsteerable, then
$K^{\prime}\Gamma_{\rho}(K^{\prime})^{\rm T}+M^{\prime}$
$=$ $A(K\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\Gamma_{\rho}\Sigma_{m+n}E\Sigma_{m+n})K^{\rm T}A^{\rm T}$
$+AMA^{\rm T}+Y$
$=$ $A[K(0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)})K^{\rm T}+M]A^{\rm T}+Y$
$=$ $A(0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)})A^{\rm T}+Y$
$\geq$ $\pm i(0_{2m}\oplus{\Omega_{n}}),$
where
$0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)}=\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\Gamma_{\rho}\Sigma_{m+n}E\Sigma_{m+n},$
$0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)}=K\left(0_{2m}\oplus\sigma_{\rho}^{(1)}+Q_{\rho}^{(1)}\right)K^{\rm T}+M,$
the second equation dues to the maximal unsteerability of $\chi_{1}$, the third equation is because $\phi$ is maximal unsteerable, and the last inequality is as $\chi_{2}$ is maximal unsteerable. It follows that $\Phi\left(\phi\right)\in{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$.
On the other hand, suppose that the matrices $A,E$ and $Y$ satisfy
$Y+\left(0_{2m}\oplus i\Omega_{n}\right)-A\left(0_{2m}\oplus i\Omega_{n}\right){A^{\rm T}}\geq 0$ (18)
and
$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}\geq 0.$ (19)
Noting that $\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}=-\left(0_{2m}\oplus i\Omega_{n}\right)$, Ineq.(19) implies
$\left(0_{2m}\oplus i\Omega_{n}\right)$ (20)
$-\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}E\Sigma_{m+n}\geq 0.$
Comparing Ineqs.(18), (20) and (9) gives $\chi_{1},{\chi_{2}}\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$. In this case, for any $\phi\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$, $K^{\prime}$ and $M^{\prime}$ satisfies
$M^{\prime}+({0_{2m}}\oplus i{\Omega_{n}})-K^{\prime}\left({0_{2m}}\oplus i{\Omega_{n}}\right){K^{\prime\rm T}}$
$=$ ${A}MA^{\rm T}+Y+(0_{2m}\oplus i\Omega_{n})$
$-{A}{K}\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}(0_{2m}\oplus i{\Omega_{n}})\Sigma_{m+n}E\Sigma_{m+n}{{K^{\rm T}}}{{A}^{\rm T}}$
$\geq$ $AMA^{\rm T}+Y-AK(0_{2m}\oplus i\Omega_{n})K^{\rm T}A^{\rm T}+(0_{2m}\oplus i\Omega_{n})$
$=$ $A[M+(0_{2m}\oplus i\Omega_{n})-K(0_{2m}\oplus i\Omega_{n})K^{\rm T}]A^{\rm T}$
$+Y+(0_{2m}\oplus i\Omega_{n})-A(0_{2m}\oplus i\Omega_{n})A^{\rm T}$
$\geq$ $Y+(0_{2m}\oplus i\Omega_{n})-A(0_{2m}\oplus i\Omega_{n})A^{\rm T}\geq 0.$
Hence $\Phi\left(\phi\right)={\chi_{2}}\circ\phi\circ{\chi_{1}}\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
Based on the above discussions and Theorem 8, we can give sufficient conditions of Gaussian superchannels becoming maximal unsteerable and unsteerable, respectively.
Theorem 9.
Assume that $\Phi=\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian superchannel. If either
(1) there exist maximal Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},0,0)$, $\chi_{2}=\chi_{2}(A,Y,\nu)\in\mathcal{GC}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}(H_{A}\otimes H_{B})$, or
(2) the matrices $A,E$ and $Y$ satisfy
$\begin{array}[]{rl}&\mathbf{w}^{{\dagger}}Y\mathbf{w}+|\mathbf{w}^{{\dagger}}A\left(0_{2m}\oplus\Omega_{n}\right)A^{\rm T}\mathbf{w}|\\ \geq&|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|\end{array}$
and
$\begin{array}[]{rl}&|\mathbf{w}^{{\dagger}}\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\left(0_{2m}\oplus\Omega_{n}\right)\Sigma_{m+n}E\Sigma_{m+n}\mathbf{w}|\\ \geq&|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|\end{array}$
for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, then $\Phi\in{\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$.
If either
(3) there exist Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},0,0)$, $\chi_{2}=\chi_{2}(A,Y,\nu)\in\mathcal{GC}_{\mathcal{US}}(H_{A}\otimes H_{B})$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}(H_{A}\otimes H_{B})$, or
(4) the matrices $A,E$ and $Y$ satisfy Ineqs.(18)-(19), then $\Phi\in{\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
We remark here that Ineq.(19) is equivalent to the following equation
$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}=0.$
Notice that the converse of Theorem 9 may not be true. In fact, if $\chi_{2}$} is a steering-annihilating Gaussian channel, by Proposition 3, $\Phi\left(A,E,Y,\nu\right)$ always is a maximal Gaussian unsteerable superchannel, regardless of the property of $\chi_{1}$.
In addition, there exists maximal Gaussian unsteerable superchannels which are not Gaussian unsteerable superchannels. For example, take a $(2+2)$-mode Gaussian superchannel $\Phi=\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}(H_{A}\otimes H_{B})$ with
[…]
and
[…]
By calculations, we find that $A,E,Y$ satisfy
${w^{\dagger}}Yw+\left|{{w^{\dagger}}A\left({{0_{A}}\oplus{\Omega_{B}}}\right){A^{T}}w}\right|\geq\left|{{w^{\dagger}}\left({{0_{A}}\oplus{\Omega_{B}}}\right)w}\right|$}
for any $w\in\mathbb{C}^{4}$, but do not satisfy
$Y-i\left({{0_{A}}\oplus{\Omega_{B}}}\right)+A\left({{0_{A}}\oplus i{\Omega_{B}}}\right){A^{T}}\geq 0.$}
That implies that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$, but $\Phi\left(A,E,Y,\nu\right)\notin{\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
V Conclusion
As a core component of CV systems, the research on Gaussian channels holds indispensable theoretical value and practical significance. Gaussian channels not only serve as a typical and tractable research vehicle for quantum resource theory, acting as an ideal model to analyze core quantum resources such as coherence, entanglement, and EPR steering, but also deepen the understanding of fundamental physical issues including quantum system symmetry and environmental decoherence, thereby improving the axiomatic framework and mathematical formulation of quantum information theory. As one of quantum resources, EPR steering is a unique quantum resource situated between quantum entanglement and Bell nonlocality, which play significant roles in various quantum protocols, secure communication and other fields.
In this work, we investigate several classes of Gaussian channels associated with EPR steering: Gaussian steering-annihilating channels (which completely eliminate steering), Gaussian steering-breaking channels (which locally disrupt steering), and maximal Gaussian unsteerable channels. We derive the necessary and sufficient conditions for a Gaussian channel to belong to each of these classes, respectively, and clarify the relationships among them. Notably, there exist channels that are simultaneously steering-annihilating and steering-breaking. In addition, from the perspective of quantum resource theory, we discuss the structure of Gaussian unsteerable superchannels as free operations, and establish the necessary and sufficient conditions for Gaussian channels to be maximal unsteerable. These results provide deeper insights into quantum channels in CV systems. Future work may focus on constructing a resource theory for quantifying the steering capability of Gaussian channels in CV systems, as well as exploring other quantum resources—such as entanglement—for bosonic Gaussian channels and non-Gaussian channels.
VI Acknowledments
This work is partially supported by National Natural Science Foundation of China (11901421, 12571138, 12171290) and the special fund for Science and Technology Innovation Teams of Shanxi Province (202304051001035).

### v2

@@
 Abstract
-Quantum steering is a crucial quantum resource that lies intermediate between entanglement and Bell nonlocality.
+EPR steering is a crucial quantum resource that lies intermediate between entanglement and Bell nonlocality.
 Gaussian channels, meanwhile, play a foundational role in diverse quantum protocols, secure communication, and related fields.
-In this paper, we focus on several classes of Gaussian channels associated with quantum steering: Gaussian steering-annihilating channels, Gaussian steering-breaking channels, Gaussian unsteerable channels, and maximal Gaussian unsteerable channels.
+In this paper, we focus on several classes of Gaussian channels associated with EPR steering: Gaussian steering-annihilating channels, Gaussian steering-breaking channels, Gaussian unsteerable channels, and maximal Gaussian unsteerable channels.
 We give the concepts of these channels, derive the necessary and sufficient conditions for a Gaussian channel to belong to each class, and explore the intrinsic relationships among them.
@@
 […]
-Gaussian states are a special class of quantum states in continuous-variable (CV) systems, playing a pivotal role in quantum optics and quantum information theory (SLP ; WPG1 ; S1 ).
-Over the past few years, the EPR steering criteria and measures for Gaussian states have garnered considerable attention from researchers (see, e.g., JiL ; KSA ; KLRA ; ZYL ; TGRF ; MKIB ; BJC ; YHQ and the references therein).
+Gaussian states are a special class of quantum states in continuous-variable (CV) quantum systems, playing a pivotal role in quantum optics and quantum information theory SLP ; WPG1 ; S1 .
+Over the past few years, the EPR steering criteria and measures for Gaussian states have garnered considerable attention from researchers (see JiL ; KSA ; KLRA ; ZYL ; TGRF ; MKIB ; BJC ; YHQ and the references therein).
 Notably, quantum information processing inevitably involves quantum channels.
-As a distinctive class of quantum channels, Gaussian quantum channels not only furnish a core theoretical framework for elucidating the inherent physical limitations of quantum communication and quantum computing, but also directly underpin the translation of quantum technologies from theoretical concepts to practical applications (BSS ; MDG ; CEGH ; CD ; NGA ; LD ; PMGH ).
+As a distinctive class of quantum channels, Gaussian quantum channels not only furnish a core theoretical framework for elucidating the inherent physical limitations of quantum communication and quantum computing, but also directly underpin the translation of quantum technologies from theoretical concepts to practical applications BSS ; MDG ; CEGH ; CD ; NGA ; LD ; PMGH .
 They thus hold irreplaceable significance in quantum systems, especially in optical quantum systems.
-One primary objective of this paper is to investigate three types of special Gaussian quantum channels with respect to EPR steering: Gaussian steering-annihilating channels that completely eliminate steering; Gaussian steering-breaking channels that locally disrupt steering; and maximal Gaussian unsteerable channels that map Gaussian unsteerable states into Gaussian unsteerable states.
+In quantum information processing, the storage and transmission of quantum states are crucial but inevitably affected by the environment.
+Thus, researching which types of environmental influences should be avoided and which are acceptable is of critical importance.
+In other words, analyzing the structure of quantum channels is essential.
+One promising research direction focuses on the dynamics of quantum resources, such as quantum entanglement, EPR steering and non-locality, under local quantum channels.
+Horodecki, Shor and Ruskai HSR studied entanglement-breaking channels in bipartite discrete-variable quantum systems which completely disentangle the subsystem they act on from the rest of the system, and proved that the Holevo capacity of entanglement-breaking channels is additive.
+This work provides theoretical support for optimizing classical information transmission over noisy channels in quantum communication.
+Holevo HEB discussed the condition any mode Gaussian quantum channels becoming entanglement-breaking, and using this condition to study several additivity conjectures of capacity for one mode Gaussian channels.
+The results in KGEB indicate that Gaussian entanglement-breaking channels and entanglement-annihilating channels serve as essential benchmarks for the infeasibility assessment of entanglement preservation, and act as core tools for both theoretical analysis and practical guidance in this field.
+Based on Gaussian entanglement-breaking channels, passive eavesdropping-immune CV quantum key distribution protocols can be constructed.
+Pereg P11 investigated the communication performance of entanglement-breaking channels with unreliable entanglement assistance, revealing the mechanism by which initial entanglement enhances the signal-to-noise ratio through classical correlations.
+Hence the study on (Gaussian) entanglement-breaking channels is of great practical significance.
+Note that EPR steering is another important quantum resource.
+Inspired by the above works, the primary objective of this paper is to investigate some types of special Gaussian quantum channels with respect to EPR steering: Gaussian steering-breaking channels that locally disrupt steering; Gaussian steering-annihilating channels that completely eliminate steering; and maximal Gaussian unsteerable channels that map Gaussian unsteerable states into Gaussian unsteerable states, so as to exploit the ability of Gaussian channels to create or destroy steering and  lay a solid theoretical foundation for the resource theory of steering for Gaussian channels.
 Regarding the correlation measures and resource theories of quantum channels, substantial research efforts have been devoted.
@@
 $\begin{array}[]{rl}{\mathbf{d}}=&(\langle\hat{R}_{1}\rangle,\langle\hat{R}_{2}\rangle,\ldots,\langle\hat{R}_{2N}\rangle)^{\rm T}\\ =&({\rm tr}(\rho R_{1}),{\rm tr}(\rho R_{2}),\ldots,{\rm tr}(\rho R_{2N}))^{\rm T}\in{\mathbb{R}}^{2N}\end{array}$
-is called the mean or the displacement vector of $\rho$ and $\Gamma=(\gamma_{kl})\in\mathcal{M}_{2N}(\mathbb{R})$ is called the covariance matrix (CM) of $\rho$ defined by $\gamma_{kl}={\rm tr}[\rho(\Delta\hat{R}_{k}\Delta\hat{R}_{l}+\Delta\hat{R}_{l}\Delta\hat{R}_{k})]$ with $\Delta\hat{R}_{k}=\hat{R}_{k}-\langle\hat{R}_{k}\rangle$ (SLP ).
+is called the mean or the displacement vector of $\rho$ and $\Gamma=(\gamma_{kl})\in\mathcal{M}_{2N}(\mathbb{R})$ is called the covariance matrix (CM) of $\rho$ defined by $\gamma_{kl}={\rm tr}[\rho(\Delta\hat{R}_{k}\Delta\hat{R}_{l}+\Delta\hat{R}_{l}\Delta\hat{R}_{k})]$ with $\Delta\hat{R}_{k}=\hat{R}_{k}-\langle\hat{R}_{k}\rangle$ SLP .
 Here, $\mathcal{M}_{d}(\mathbb{R})$ stands for the algebra of all $d\times d$ matrices over the real field $\mathbb{R}$.
@@
 Note that $\Gamma$ is real symmetric and satisfies the condition
-$\Gamma+i\Omega_{N}\geq 0,\ \ {\rm where}\ \ \Omega_{N}=\oplus_{k=1}^{N}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
+$\Gamma+i\Omega_{N}\geq 0,$
+where
+$\Omega_{N}=\underbrace{\Omega\oplus\cdots\oplus\Omega}_{N}\ \ {\rm with}\ \ \Omega=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$.
 Now, divide the $N$-mode CV system into $m$-mode CV subsystem A and $n$-mode CV subsystem B, with state space $H=H_{A}\otimes H_{B}$ and $N=m+n$.
@@
 where $a,b\geq 1,\ ab-c^{2}\geq 1,\ ab-d^{2}\geq 1$. For more details about Gaussian states, see WPG1 ; S1 .
-In the rest of this paper, if there is no explanation, we always assume that $H$, $H_{A}$ and $H_{B}$ are infinite-dimensional separable complex Hilbert spaces with $H=H_{A}\otimes H_{B}$.
 II.2 EPR steering
-A measurement assemblage ${\mathcal{MA}}=\{M_{a|x}\}_{a,x}$ is a collection of positive operators $M_{a|x}\geq 0$ satisfying $\sum_{a}M_{a|x}=I$ for each $x$.
-Such a collection represents one positive-operator-valued measurement (POVM), describing a general quantum measurement, for each $x$.
-In a (bipartite) EPR steering scenario, one party performs measurements on a shared state $\rho_{AB}\in{\mathcal{S}}(H_{A}\otimes H_{B})$, which steers the quantum state of the other particle.
-If Alice performs a set of measurements $\{M^{A}_{a|x}\}_{a,x}$, then the collection of sub-normalized of Bob is an assemblage $\{\rho^{B}_{a|x}\}_{a,x}$ with
+In a bipartite EPR steering scenario, Alice and Bob share a bipartite state $\rho_{AB}\in{\mathcal{S}}(H_{A}\otimes H_{B})$, and Alice performs positive-operator-valued measurements (POVMs) on her subsystem to steer $\rho_{AB}$ on Bob’s side.
+If Alice performs a set of POVMs ${\mathcal{MA}}=\{M_{a|x}\}_{a,x}$ (that is, $M_{a|x}\geq 0$ with $\sum_{a}M_{a|x}=I$ for each $x$), then the assemblage of sub-normalized “conditional states” of the subsystem state $\rho_{B}={\rm Tr}_{A}(\rho_{AB})$ is $\{\rho^{B}_{a|x}\}_{a,x}$, where
 $\rho^{B}_{a|x}={\rm Tr}_{A}((M^{A}_{a|x}\otimes I_{B}){\rho}_{AB}).$
-If every assemblage on Bob $\{\rho^{B}_{a|x}\}_{a,x}$ can be explained by a local hidden state (LHS) model:
+$\rho_{AB}$ is said to be unsteerable from A to B if every assemblage $\{\rho^{B}_{a|x}\}_{a,x}$ on Bob’s side can be explained by a local hidden state (LHS) model as follows
 $\rho^{B}_{a|x}=\sum_{\lambda}p_{\lambda}p(a|x,\lambda)\sigma_{\lambda},$
-where $\lambda$ is a hidden variable, distributed according to $p_{\lambda}$, $\sigma_{\lambda}$ are “hidden states” of Bob, and $p(a|x,\lambda)$ are local “response functions” of Alice, then we say that $\rho_{AB}$ has a LHS form, or does not demonstrate steering (WJ ).
-Otherwise, if there exist measurements such that $\rho^{B}_{a|x}$ does not admit such LHS decomposition, then $\rho_{AB}$ is called steerable from A to B.
-Symmetrically, we can define the steerability of $\rho_{AB}$ from B to A.
-Here, we point out that, if there is no special illustration, we focus on the steerability from A to B in this paper.
+where $\lambda$ is a hidden variable, $p_{\lambda}$ is a distribution in $\lambda$, $p(a|x,\lambda)$ are local “response functions” of Alice, and $\sigma_{\lambda}$ are “hidden states” of Bob.
+Otherwise, $\rho_{AB}$ is called steerable (from A to B) WJ .
+Symmetrically, we can define the steerability of $\rho_{AB}$ (from B to A).
 In CV systems, Gaussian POVM (GPOVM) plays an important role. Recall that an $N$-mode GPOVM $\Pi=\{\Pi(\alpha)\}$ is defined as
 $\Pi(\alpha)=\frac{1}{\pi^{N}}D(\alpha)\varpi D^{\dagger}(\alpha),$
-where $\alpha=(\alpha_{1},\cdots,\alpha_{N})^{\rm T}\in{\mathbb{C}^{N}}$, $D(\alpha)={\rm exp}[\sum_{j=1}^{N}(\alpha_{j}\hat{a}_{j}^{\dagger}-\alpha_{j}^{*}\hat{a}_{j})]$ is the $N$-mode Weyl displacement operator and $\varpi$ is a zero mean $N$-mode Gaussian state, which is called the seed state of $\Pi$ (GC ).
+where $\alpha=(\alpha_{1},\cdots,\alpha_{N})^{\rm T}\in{\mathbb{C}^{N}}$, $D(\alpha)={\rm exp}[\sum_{j=1}^{N}(\alpha_{j}\hat{a}_{j}^{\dagger}-\alpha_{j}^{*}\hat{a}_{j})]$ is the $N$-mode Weyl displacement operator and $\varpi$ is a zero mean $N$-mode Gaussian state, which is called the seed state of $\Pi$ GC .
 For any bipartite Gaussian state, the authors in WJ derived a linear matrix inequality steering criterion via GPOVMs.
 Assume that $\rho\in{\mathcal{S}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3).
-As demonstrated in WJ , $\rho$ is unsteerable by the subsystem A’s all GPOVMs if and only if
-$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0,\ \ {\rm where}\ \ \Omega_{n}=\oplus_{k=1}^{n}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$. (5)
-Obviously,
-$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\Leftrightarrow\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n}).$
-Denote respectively by ${\mathcal{GS}}(H)$ and ${\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ the set of all Gaussian states on $H$ and the set of all Gaussian unsteerable states from A to B on $H_{A}\otimes H_{B}$, that is,
-$\mathcal{GS}(H)=\{\rho\in{\mathcal{S}}(H):\ \rho\ \mbox{is a Gaussian state}\}$
+As demonstrated in WJ , $\rho$ is unsteerable (from A to B)by the subsystem A’s all GPOVMs if and only if
+$\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0;$ (5)
+and $\rho$ is unsteerable by the subsystem B’s all GPOVMs if and only if
+$\Gamma_{\rho}+i\Omega_{m}\oplus 0_{2n}\geq 0.$
+Note that any positive semidefnite matrix and its transpose matrix have the same eigenvalues. So the condition (5) is equivalent to
+$(\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n})^{\rm T}\geq 0,$
+that is,
+$\Gamma_{\rho}-0_{2m}\oplus i\Omega_{n}\geq 0.$
+Hence
+$\begin{array}[]{rl}&\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\\ \Leftrightarrow&\Gamma_{\rho}\geq\pm(0_{2m}\oplus i\Omega_{n})=\pm i(0_{2m}\oplus\Omega_{n}).\end{array}$
+Denote respectively by ${\mathcal{GS}}^{(N)}$ and $\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$} the set of all $N$-mode Gaussian states, the set of all $(m+n)$-mode Gaussian states that are Gaussian unsteerable (from A to B) by using Alice’s Gaussian measurements.
+Particularly, if $N=m+n$, denote ${\mathcal{GS}}^{(N)}={\mathcal{GS}}^{(m,n)}.$ That is,
+${\mathcal{GS}}^{(N)}=\{\mbox{all N-mode Gaussian states}\}$
 and
-$\begin{array}[]{rl}&{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =&\{\rho\in{\mathcal{GS}}(H_{A}\otimes H_{B}):\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\}.\end{array}$
+$\begin{array}[]{rl}&{\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}}\\ =&\{\rho\in{\mathcal{GS}}^{(m,n)}:\Gamma_{\rho}+0_{2m}\oplus i\Omega_{n}\geq 0\}\\ =&\{\rho\in{\mathcal{GS}}^{(m,n)}:\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n})\}.\end{array}$
 II.3 Gaussian unsteerable channels
-Recall that a Gaussian channel is a quantum channel which transforms Gaussian states into Gaussian states (PMGH ; CEGH ).
-An $N$-mode Gaussian channel $\phi$ on $H$ can be described by $\phi=\phi\left({K,M,\mathbf{d}}\right)$, which acts on $\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in\mathcal{GS}(H)$ as
+Recall that a Gaussian channel is a quantum channel which transforms any Gaussian states into Gaussian statesPMGH ; CEGH .
+An $N$-mode Gaussian channel $\phi$ can be described by $\phi=\phi\left({K,M,\mathbf{d}}\right)$, which acts on $\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in\mathcal{GS}(N)$ as
 $\mathbf{d}_{\rho}\mapsto K\mathbf{d}_{\rho}+\mathbf{d},\ \ \Gamma_{\rho}\mapsto K\Gamma_{\rho}K^{\rm T}+M,$ (6)
-$\mathbf{d}\in\mathbb{R}^{2N}$ is a column displacement vector, $K,M\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $M=M^{\rm T}$ and the completely positive condition
+where $\mathbf{d}\in\mathbb{R}^{2N}$ is a column displacement vector, $K,M\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $M=M^{\rm T}$ and the completely positive condition
 $M+i\Omega_{N}-iK\Omega_{N}{K^{\rm T}}\geq 0.$ (7)
-Denote by ${\mathcal{GC}}(H)$ the set of all Gaussian channels on $H$. Here, we give three known Gaussian channels which are used frequently.
-Attenuator channels. A single-mode attenuator channel $\phi_{\theta}^{n_{th}}(K,M,\mathbf{d})$ is a deterministic Gaussian channel with
+Denote by ${\mathcal{GC}}^{(N)}$ the set of all $N$-mode Gaussian channels:
+${\mathcal{GC}}^{(N)}=\{\mbox{all N-mode Gaussian quantum channels}\}.$
+Particularly, if $N=m+n$, write ${\mathcal{GC}}^{(N)}={\mathcal{GC}}^{(m,n)}$.
+Here, we give three known Gaussian channels which are used frequently.
+Attenuator channels. A single-mode attenuator channel $\phi_{\theta}^{n_{th}}(K,M,\mathbf{d})$ is a deterministic Gaussian channel S1 with
 […]
-where $\theta\in[0,2\pi]$ and thermal noise $n_{th}\geq 1$. Particularly, if $n_{th}=1$, then the channel is called a pure lossy channel.
+where $\theta\in[0,2\pi]$, the thermal noise $n_{th}=2N_{0}+1\geq 1$ and $N_{0}$ is mean number of thermal occupation. Particularly, if $n_{th}=1$, then the channel is called a pure lossy channel.
 Constant channels.
-A Gaussian channel $\Theta(K,M,\mathbf{d})\in{\mathcal{GC}}(H)$ is called a constant channel if there exists some $\rho_{0}(\Gamma_{0},{\mathbf{d}}_{0})\in\mathcal{GS}(H)$ such that $\Theta(\rho)=\rho_{0}$ for all $\rho\in\mathcal{GS}(H)$.
+An $N$-mode Gaussian channel $\Theta(K,M,\mathbf{d})$ is called a constant channel if there exists some $N$-mode Gaussian state $\rho_{0}(\Gamma_{0},{\mathbf{d}}_{0})\in\mathcal{GS}^{(N)}$ such that $\Theta(\rho)=\rho_{0}$ for all $N$-mode Gaussian states $\rho\in\mathcal{GS}^{(N)}$.
 In this case, $\Theta(K,M,\mathbf{d})$ can be represented as $\Theta(K,M,\mathbf{d})=\Theta(0,\Gamma_{0},\mathbf{\mathbf{d}}_{0})$.
 Identity Gaussian channel.
-A Gaussian channel $\psi(K,M,\mathbf{d})\in{\mathcal{GC}}(H)$ is called an identity channel if $\psi(\rho)=\rho$ holds for all $\rho\in\mathcal{GS}(H)$.
+An $N$-mode Gaussian channel $\psi(K,M,\mathbf{d})$ is called an identity channel if $\psi(\rho)=\rho$ holds for all $N$-mode Gaussian states $\rho\in\mathcal{GS}^{(N)}$.
 In this case, $K=I$, $M=0$ and $\mathbf{d}=0$.
 In YHQ , the authors gave the definition of Gaussian unsteerable channels.
-Any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}(H_{A}\otimes H_{B})$ is called Gaussian unsteerable (from A to B) if $K$ and $M$ satisfy the following relation
+Recall that any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)$ is called Gaussian unsteerable from A to B if $K$ and $M$ satisfy the following relation
 $M+({0_{2m}}\oplus i{\Omega_{n}})-K\left({0_{2m}}\oplus i{\Omega_{n}}\right){K^{\rm T}}\geq 0;$ (9)
-and is called maximal Gaussian unsteerable (from A to B) if $\phi({\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})$. Let
-$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =&\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}(H_{A}\otimes H_{B}):\\ &M+(0_{2m}\oplus i{\Omega_{n}})-K\left(0_{2m}\oplus i{\Omega_{n}}\right){K^{\rm T}}\geq 0\}\end{array}$
+and is called maximal Gaussian unsteerable from A to B if $\phi$ maps all $(m+n)$-mode Gaussian unsteerable states from A to B into $(m+n)$-mode Gaussian unsteerable states from A to B.
+Let ${\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ and ${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ stand for the set of all $(m+n)$-mode Gaussian unsteerable (from A to B) channels and the set of all $(m+n)$-mode maximal Gaussian unsteerable (from A to B) channels, respectively, that is,
+${\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}=\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}^{(m,n)}:$
+$M+(0_{2m}\oplus i{\Omega_{n}})-K\left(0_{2m}\oplus i{\Omega_{n}}\right){K^{\rm T}}\geq 0\},$
 and
-$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\\ =&\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}(H_{A}\otimes H_{B}):\\ &\phi({\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H_{A}\otimes H_{B})\}.\end{array}$
-It is shown in YHQ that the set ${\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ is a proper (but) large subset of ${\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$.
-Obviously, the $(m+n)$-mode identity Gaussian channel $\phi(I,0,0)\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
+${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}=\{\phi(K,M,\mathbf{d})\in{\mathcal{GC}}^{(m,n)}:$
+$\phi(\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)})\subseteq\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}\}.$
+It is shown in YHQ that the set ${\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$is a proper (but) large subset of ${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$.
+Obviously, the $(m+n)$-mode identity Gaussian channel $\phi(I,0,0)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.
 II.4 Gaussian superchannels
-Recall that a superchannel is a completely positive linear map transforming any quantum channels into quantum channels; and a Gaussian superchannel is a superchannel transforming any Gaussian channels into Gaussian channels (X11 ).
-Write
-${\mathcal{SGC}}(H)=\{\mbox{all Gaussian superchannels on }H\}.$
-Xu X11 gave two different representations of Gaussian superchannels.
-Theorem 1.
-(X11 ) Any $N$-mode Gaussian superchannel $\Phi\in{\mathcal{SGC}}(H)$ can be represented by $\Phi\left({A,E,Y,\nu}\right)$, where $\nu\in{\mathbb{R}}^{2N}$, $A,E,Y\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $Y=Y^{\rm T}$, $EE^{\rm T}={I_{2N}}$ and
-$Y+i\Omega_{N}-iA\Omega_{N}A^{\rm T}\geq 0,\ \ i\Omega_{N}-iE\Omega_{N}E^{\rm T}\geq 0.$ (10)
-Moreover, for any $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}(H)$, we have
-(1) $\Phi\left({\phi\left({K,M,\mathbf{d}}\right)}\right)=\phi^{\prime}\left({K^{\prime},M^{\prime},\mathbf{d}^{\prime}}\right)$ with
-$K^{\prime}=AK\Sigma_{N}E^{\rm T}\Sigma_{N},\ M^{\prime}=AMA^{\rm T}+Y,\ \mathbf{d}^{\prime}=A\mathbf{d}+\nu,$
-[…]
-(2) $\Phi\left(\phi\left({K,M,\mathbf{d}}\right)\right)={\phi_{2}}\circ\phi\circ{\phi_{1}}$ with some ${\phi_{1}}\left(K_{1},M_{1},\mathbf{d}_{1}\right)$ and ${\phi_{2}}\left(K_{2},M_{2},\mathbf{d}_{2}\right)\in{\mathcal{GC}}(H)$, where $K_{1}=\Sigma_{N}E^{\rm T}\Sigma_{N},$ $M_{1}=0,$ ${\mathbf{d}_{1}}=0,$ $K_{2}=A,$ $M_{2}=Y$, ${\mathbf{d}_{2}}=\upsilon.$
+Recall that a superchannel is a completely positive linear map transforming any quantum channels into quantum channels; and a Gaussian superchannel is a superchannel transforming any Gaussian channels into Gaussian channels X11 .
+Denote by
+${\mathcal{GSC}}^{(N)}=\{\mbox{all N-mode Gaussian superchannels}\}.$
+It is shown in X11 that any $N$-mode Gaussian superchannel $\Phi\in{\mathcal{GSC}}^{(N)}$ can be described by $\Phi\left({A,E,Y,\nu}\right)$ as follows: for any $N$-mode Gaussian channel $\phi\left({K,M,\mathbf{d}}\right)$, $\Phi\left({\phi\left({K,M,\mathbf{d}}\right)}\right)=\phi^{\prime}\left({K^{\prime},M^{\prime},\mathbf{d}^{\prime}}\right)$ with
+$K^{\prime}=AK\Sigma_{N}E^{\rm T}\Sigma_{N},$
+$M^{\prime}=AMA^{\rm T}+Y,$
+$\mathbf{d}^{\prime}=A\mathbf{d}+\nu,$
+where $A,E,Y\in\mathcal{M}_{2N}(\mathbb{R})$, $Y=Y^{\rm T}$, $EE^{\rm T}={I_{2N}}$, $\nu\in{\mathbb{R}}^{2N}$, $\Sigma_{N}=\underbrace{\Sigma\oplus\cdots\oplus\Sigma}_{N}$ with $\Sigma={\begin{pmatrix}1&0\\ 0&{-1}\end{pmatrix}}$, and
+$Y+i\Omega_{N}-iA\Omega_{N}A^{\rm T}\geq 0,$
+$i\Omega_{N}-iE\Omega_{N}E^{\rm T}\geq 0.$ (10)
+In addition, the author X11 also proved that any $N$-mode Gaussian superchannel $\Phi\left({A,E,Y,\nu}\right)$ can be described in terms of compositions of Gaussian channels, that is, for any $N$-mode Gaussian channel $\phi\left({K,M,\mathbf{d}}\right)$, there exist two $N$-mode Gaussian channels ${\phi_{1}}\left(K_{1},M_{1},\mathbf{d}_{1}\right)$ and ${\phi_{2}}\left(K_{2},M_{2},\mathbf{d}_{2}\right)$ such that $\Phi(\phi)={\phi_{2}}\circ\phi\circ{\phi_{1}}$.
+Particularly, one such representation is
+$\begin{cases}K_{1}=\Sigma_{N}E^{\rm T}\Sigma_{N},\ \ M_{1}=0,\ \ {\mathbf{d}_{1}}=0;\\ K_{2}=A,\ \ M_{2}=Y,\ \ {\mathbf{d}_{2}}=\upsilon.\end{cases}$
 III Gaussian steering-annihilating and Gaussian steering-breaking channels
@@
 Definition 1.
-Assume that $H,K$ are any infinite-dimensional separable complex Hilbert spaces.
-We say that any Gaussian channel $\phi\in{\mathcal{GC}}(H\otimes K)$ is Gaussian steering-annihilating if it sends all Gaussian states into Gaussian unsteerable states, that is, if $\phi(\mathcal{GS}(H\otimes K))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H\otimes K)$; and any Gaussian channel $\psi\in\mathcal{GC}(H)$ is Gaussian steering-breaking if $(\psi\otimes I_{K})(\rho)$ is always unsteerable for all Gaussian states $\rho\in{\mathcal{GS}}(H\otimes K)$, that is, if $(\psi\otimes I_{K})(\mathcal{GS}(H\otimes K))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H\otimes K)$.
-Let
-$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\\ =\{&\mbox{all Gaussian steering-annihilating }\\ &\mbox{channels on }H_{A}\otimes H_{B}\}\end{array}$
-and
-$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{SB}}(H)\\ =&\{\mbox{ all Gaussian steering-breaking channels on }H\}.\end{array}$
-By Definition 2, the following useful property is obvious.
+Assume that $\phi$ is any $(m+n)$-mode bipartite Gaussian channel.
+We say that $\phi$ is steering-annihilating if it sends all $(m+n)$-mode Gaussian states into either $(m+n)$-mode Gaussian unsteerable states from A to B or $(m+n)$-mode Gaussian unsteerable states from B to A.
+Definition 2.
+Assume that $\psi$ is any $m$-mode Gaussian channel.
+We say that $\psi$ is steering-breaking if for any $(m+n)$-mode Gaussian state $\rho$, $(\psi\otimes I_{K})(\rho)$ is always $(m+n)$-mode Gaussian unsteerable from A to B or from B to A, where $K$ denotes arbitrary $n$-mode ancillary CV system.
+For the convinience, denote by ${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$, ${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}(A\rightarrow B)}$ the set of all $(m+n)$-mode Gaussian steering-annihilating channels and the set of all $(m+n)$-mode Gaussian steering-annihilating channels sending all $(m+n)$-mode Gaussian states into $(m+n)$-mode Gaussian unsteerable states from A to B, respectively.
+Denote by ${\mathcal{GC}}^{(N)}_{\mathcal{SB}}$, ${\mathcal{GC}}^{(N)}_{\mathcal{SB}(A\rightarrow B)}$ the set of all $N$-mode Gaussian steering-breaking channels and the set of all $N$-mode Gaussian steering-breaking channels from A to B, respectively.
+By Definitions 1 and 2, the following useful property is obvious.
 Proposition 3.
-If $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$, and $\psi\in{\mathcal{GC}}(H_{A}\otimes H_{B})$, then $\phi\circ\psi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$; if $\phi\in{\mathcal{GC}}_{\mathcal{SB}}(H)$ and $\psi\in{\mathcal{GC}}(H)$, then both $\phi\circ\psi$ and $\psi\circ\phi$ belong to ${\mathcal{GC}}_{\mathcal{SB}}(H)$.
+If $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$, and $\psi\in{\mathcal{GC}}^{(m,n)}$, then $\phi\circ\psi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$; if $\phi\in{\mathcal{GC}}^{(N)}_{\mathcal{SB}}$ and $\psi\in{\mathcal{GC}}^{(N)}$, then both $\phi\circ\psi$ and $\psi\circ\phi$ belong to ${\mathcal{GC}}^{(N)}_{\mathcal{SB}}$.
 Next, we first give a sufficient condition for Gaussian channels being Gaussian steering-annihilating.
 Theorem 4.
-Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any $(m+n)$-mode channel. If $\phi$ satisfies the condition
-$M+({0_{2m}}\oplus i{\Omega_{n}})-iK\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\geq 0,$ (11)
-then $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$.
+Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian channel. If $\phi$ satisfies one of the following conditions:
+(1) $M+({0_{2m}}\oplus i{\Omega_{n}})-iK\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\geq 0,$
+(2) $M+(i{\Omega_{m}}\oplus{0_{2n}})-iK\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\geq 0,$
+then $\phi$ is steering-annihilating, that is, $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$.
 Proof.
-Assume that $\rho\in{\mathcal{GS}}\left(H_{A}\otimes H_{B}\right)$ is any Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3).
+Assume that $\rho\in{\mathcal{GS}}^{(m,n)}$ is any Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3).
 Then $\phi(\rho)$ has the CM $\Gamma_{\phi(\rho)}=K\Gamma_{\rho}{K^{\rm T}}+M$.
-As $\Gamma_{\rho}+i\left({{\Omega_{m}}\oplus{\Omega_{n}}}\right)\geq 0$, by the assumption (11), we have
+As $\Gamma_{\rho}+i(\Omega_{m}\oplus\Omega_{n})\geq 0$, we have
+$K(\Gamma_{\rho}+i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T}\geq 0$ (11)
+for any $K\in\mathcal{M}_{2(m+n)}(\mathbb{R})$. If the condition (1) holds, combining this assumption with Ineq.(11), one gets
 $\Gamma_{\phi(\rho)}+(0_{2m}\oplus i{\Omega_{n}})$
 $=$ $K\Gamma_{\rho}{K^{\rm T}}+M+(0_{2m}\oplus i{\Omega_{n}})$
-$\geq$ $K(-i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T}+M+(0_{2m}\oplus i\Omega_{n})\geq 0.$
-It follows that $\phi\left(\rho\right)\in{\mathcal{GS}}_{\mathcal{U}S}(H_{A}\otimes H_{B})$. So $\phi$ is Gaussian steering-annihilating. ∎
-Notice that the condition (11) is only sufficient but not necessary for a channel being steering-annihilating.
-In fact, there exist Gaussian steering-annihilating channels which do not satisfy the condition (11).
+$=$ $K(\Gamma_{\rho}+i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T}$
+$+(M+(0_{2m}\oplus i\Omega_{n})-K(i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T})\geq 0.$
+It follows from the steering criterion (Ineq.(5)) that $\phi\left(\rho\right)\in\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.
+Similarly, one can show that, if the condition (2) holds, then $\phi\left(\rho\right)\in\mathcal{GS}^{(m,n)}_{\mathcal{US}(B\rightarrow A)}$. Hence $\phi$ is Gaussian steering-annihilating. ∎
+Notice that the conditions in Theorem 4 is only sufficient but not necessary for a channel being steering-annihilating.
+In fact, there exist Gaussian steering-annihilating channels which do not satisfy this condition.
 Example 5.
@@
 […]
-Then $\phi_{1}$ is steering-annihilating, but does not satisfy the condition (11).
+Then $\phi_{1}$ is steering-annihilating from A to B, but does not satisfy the condition (1) in Theorem 4.
 In fact, it is easily checked that $\phi\left({{K_{1}},{M_{1}},{\rm d_{1}}}\right)$ satisfies the condition
@@
 […]
-So $\phi$ does not satisfy the condition (11).
+So $\phi$ does not satisfy the condition (1) in Theorem 4.
 However, for any $(1+1)$-mode Gaussian state $\rho$ with the standard CM $\Gamma_{\rho}$ in Eq.(4), by a numerical calculation, one can obtain
 […]
-which implies that $\phi$ is steering-annihilating.
-While Ineq.(11) is not a necessary condition, the subsequent Ineq.(12) provides a sufficient and necessary condition for a Gaussian channel becoming steering-annihilating.
+which implies that $\phi$ is steering-annihilating (from A to B).
+While the condition (1) or (2) in Theorem 4 is not a necessary condition, the subsequent inequality in Theorem 6 provides a sufficient and necessary condition for a Gaussian channel becoming steering-annihilating.
 Theorem 6.
-Any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$ if and only if
+Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian channel.
+Then $\phi$ is Gaussian steering-annihilating, that is, $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$, if and only if
 $\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$ (12)
 $\geq$ $|\mathbf{w}^{{\dagger}}({0_{2m}}\oplus{\Omega_{n}})\mathbf{w}|$
+holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, or
+$\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$ (13)
+$\geq$ $|\mathbf{w}^{{\dagger}}({\Omega_{m}}\oplus{0_{2n}})\mathbf{w}|$
 holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
-Proof.
-Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any Gaussian channel.
-If $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$, then for any $\rho\in{\mathcal{GS}}\left(H_{A}\otimes H_{B}\right)$ with CM $\Gamma_{\rho}$, by Eqs.(5)-(6), one has
-$K\Gamma_{\rho}K^{\rm T}+M\geq\pm i(0_{2m}\oplus\Omega_{n}),$
-and so
-$\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}$
-holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
-Note that $\Gamma_{\rho}\geq\pm i(\Omega_{m}\oplus\Omega_{n})$, and $\inf\limits_{\sigma\geq\pm i\Omega_{N}}{\bf w}^{\dagger}\sigma{\bf w}=|{\bf w}^{\dagger}\Omega_{N}{\bf w}|$ for all $N$-mode quantum states $\sigma$ PMGH .
-Thus, the above inequality implies
-$|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
-for all $\mathbf{w}\in\mathbb{C}^{2(m+n)},$ that is, Ineq.(12) holds.
-On the other hand, if Ineq.(12) holds, then for any $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, one gets
-$\mathbf{w}^{{\dagger}}\Gamma_{\phi(\rho)}\mathbf{w}$ $=$ $\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
-$\geq$ $\inf_{\Gamma_{\rho}\geq\pm i\left({\Omega_{m}\oplus{\Omega_{n}}}\right)}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
-$=$ $|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}$
-$\geq$ $|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
-$\geq$ $\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.$
-So $\Gamma_{\phi(\rho)}\geq\pm i(0_{2m}\oplus\Omega_{n})$, which means that $\phi(\rho)$ is unsteerable. It follows that $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$. ∎
-Note that
-$\left|{{\varphi_{r}}}\right\rangle=\frac{1}{\cosh r}{\sum\limits_{j=0}^{\infty}{\left({\tanh r}\right)}^{j}}\left|j_{H}\right\rangle\left|j_{H}\right\rangle\in H\otimes H$
-is a Gaussian squeezed pure state, where $\{|j_{H}\rangle\}$ is the Fock basis of $H$ and $r\in\mathbb{R}$ is the squeezed parameter. For any Gaussian channel $\phi\in{\mathcal{GC}}(H)$, define
-${\rho_{\phi}}=(\phi\otimes I_{H})(\left|{{\varphi_{r}}}\right\rangle\left\langle{{\varphi_{r}}}\right|).$
-Clearly, $\rho_{\phi}\in{\mathcal{GS}}(H\otimes H)$. By the unsteerability of $\rho_{\phi}$, we can give necessary and sufficient conditions for Gaussian steering-breaking channels.
+Take any $N$-mode Gaussian channel $\phi\in{\mathcal{GC}}^{(N)}$ with the state space $H$. Note that the pure state
+$\left|{{\varphi_{r}}}\right\rangle=\frac{1}{\cosh r}{\sum\limits_{j=0}^{\infty}{\left({\tanh r}\right)}^{j}}\left|j\right\rangle\left|j\right\rangle$
+is a two-mode Gaussian squeezed pure state, where $\{|j\rangle|j\rangle\}$ is a tensor product of Fock states and $r\in\mathbb{R}$ is the squeezed parameter.
+Correspondingly, $2N$-mode Gaussian squeezed pure state can be written as $\left|{{\psi_{r}}}\right\rangle=\left|{{\varphi_{r}}}\right\rangle^{\otimes N}.$ Define
+${\rho_{\phi}}=(\phi\otimes I_{H})(\left|{{\psi_{r}}}\right\rangle\left\langle{{\psi_{r}}}\right|).$
+Clearly, $\rho_{\phi}\in{\mathcal{GS}}^{(N,N)}$ is a $(N+N)$-mode Gaussian state.
+By the unsteerability of $\rho_{\phi}$, we can give necessary and sufficient conditions for Gaussian steering-breaking channels.
 Theorem 7.
-Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H\right)$ is any N-mode Gaussian channel. Then the following statements are equivalent.
-(1) $\phi\in{\mathcal{GC}}_{\mathcal{SB}}(H)$ is Gaussian steering-breaking.
-(2) ${\rho_{\phi}}\in{\mathcal{GS}}_{\mathcal{US}}(H\otimes H)$.
+Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(N)}$ is any $N$-mode Gaussian channel. Then the following statements are equivalent.
+(1) $\phi$ is Gaussian steering-breaking, that is, $\phi\in{\mathcal{GC}}^{(N)}_{\mathcal{SB}}$.
+(2) ${\rho_{\phi}}\in{\mathcal{GS}^{(N,N)}_{\mathcal{US}(A\rightarrow B)}}$ or ${\rho_{\phi}}\in{\mathcal{GS}^{(N,N)}_{\mathcal{US}(B\rightarrow A)}}$.
 (3) The matrices $K$ and $M$ satisfy the condition
-$M-iK\Omega_{N}K^{\rm T}\geq 0.$
-Proof.
-Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H\right)$ is any $N$-mode Gaussian channel.
-$\left(1\right)\Rightarrow\left(2\right)$: By Definition 2, this is obvious.
-$\left(2\right)\Rightarrow\left(3\right)$: Assume that ${\rho_{\phi}}\in{\mathcal{GS}}_{\mathcal{US}}(H\otimes H)$. Note that the CM $\Gamma_{{\rho_{\phi}}}$ of $\rho_{\phi}$} has the form X11
-[…]
+$M-iK\Omega_{N}K^{\rm T}\geq 0$
+or
+$M+i\Omega_{N}\geq 0.$
+Proofs of Theorems 6 and 7 are given in Appendix.
+In the end of this section, we will investigate the relationship between Gaussian steering-breaking channels, Gaussian steering-annihilating channels and maximal Gaussian unsteerable channels.
+Denote by
+${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}}={\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}\bigcup{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(B\rightarrow A)}.$
+By their definitions, it is obvious that
+${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}={\mathcal{GC}}^{(m,n)}_{\mathcal{SA}(A\rightarrow B)}\bigcup{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}(B\rightarrow A)},$
+${\mathcal{GC}}^{(m,n)}_{\mathcal{SB}}={\mathcal{GC}}^{(m,n)}_{\mathcal{SB}(A\rightarrow B)}\bigcup{\mathcal{GC}}^{(m,n)}_{\mathcal{SB}(B\rightarrow A)}$
 and
-$\Gamma_{{\rho_{\phi}}}+(0_{2N}\oplus i{\Omega_{N}})\geq 0.$ (14)
-It is well known that a Hermitian matrix $W=\left(\begin{array}[]{cc}W_{11}&W_{12}\\ W_{12}^{{\dagger}}&W_{22}\end{array}\right)\geq 0$ if and only if $W_{22}\geq 0$ and $W_{11}-W_{12}W_{22}^{-1}W_{12}^{{\dagger}}\geq 0$ (ZH ).
-Then Ineq.(14) implies
-$\begin{cases}\cosh 2rI_{2N}+i{\Omega_{N}}\geq 0,\\ \begin{array}[]{rl}&\cosh 2rK{K^{\rm T}}+M\\ &-\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N}+i\Omega_{N})^{-1}\Sigma_{N}{K^{\rm T}}\geq 0.\end{array}\end{cases}$
-Note that $(T+cI)^{-1}=\frac{1}{c}(I-\frac{T}{c})+o(c^{-2})$ (S1 ), where $o(c^{-2})$ stands for the Landau little which will be neglected when taking to $c\to\infty$. So
-$0\leq\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N}+i{\Omega_{N}})^{-1}\Sigma_{N}K^{\rm T}]$
-$=$ $\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-\sinh^{2}(2r)K\Sigma_{N}[\frac{1}{\cosh 2r}(I_{2N}-\frac{i\Omega_{N}}{\cosh 2r})+o(\frac{1}{\cosh 2r})]\Sigma_{N}{K^{\rm T}}]$
-$=$ $\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-K\Sigma_{N}[\frac{\sinh^{2}(2r)}{\cosh 2r}I_{2N}-i\tanh^{2}(2r){\Omega_{N}}+\sinh^{2}(2r)o(\frac{1}{\cosh 2r})]\Sigma_{N}K^{\rm T}]=M-iK{\Omega_{N}}{K^{\rm T}}.$
-[…]
-[…]
-Since $\Gamma_{\rho}+i({\Omega_{N}}\oplus{\Omega_{N^{\prime}}})\geq 0$, a direct calculation gives
-$Y+i{\Omega_{N^{\prime}}}\geq 0$}
-and
-$X+i{\Omega_{N}}-Z\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}{Z^{\rm T}}\geq 0$}.
-This implies
-$KX{K^{\rm T}}+M-KZ{\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}}{Z^{\rm T}}{K^{\rm T}}$
-$=$ $K\left({X-Z{{\left({Y+i{\Omega_{N^{\prime}}}}\right)}^{-1}}{Z^{\rm T}}}\right){K^{\rm T}}+M$
-$\geq$ $K\left({-i{\Omega_{N}}}\right){K^{\rm T}}+M=M-iK{\Omega_{N}}{K^{\rm T}}\geq 0.$
-Also note that
-$KX{K^{\rm T}}+M\geq 0$
-by the denifition of the channel $\phi$.
-It follows that ${\Gamma_{\left({\phi\otimes I_{H^{\prime}}}\right)(\rho)}}+i(0_{N}\oplus{\Omega_{N^{\prime}}})\geq 0$.
-So $(\phi\otimes{I_{H^{\prime}}})\left(\rho\right)$ is unsteerable, and hence $\phi$ is steering-breaking.
-∎
-In the end of this section, we will investigate the relationship between Gaussian steering-breaking channels, Gaussian steering-annihilating channels and maximal Gaussian unsteerable channels.
-By their definitions, it is obvious that
-${\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\subset{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}).$
-As a consequence of Proposition 3, for any $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$ and $\psi\in{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})$, we have
-$\phi\circ\psi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\cap{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B}),$
+${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}\subset{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}}$.
+As a consequence of Proposition 3, for any $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$ and $\psi\in{\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}$, we have
+$\phi\circ\psi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}\cap{\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}$,
 which means that there are Gaussian channels which are simultaneously steering-breaking and steering-annihilating.
@@
 Obviously, by Theorem 7, $\Theta$ is steering-breaking.
-However, if $\Gamma_{0}+(0_{2m}\oplus i\Omega_{n})\ngeq 0$, $\Theta$ is neither steering-annihilating nor maximal Gaussian unsteerable.
+However, if $\Gamma_{0}+(0_{2m}\oplus i\Omega_{n})\ngeq 0$, and $\Gamma_{0}+(i\Omega_{m}\oplus 0_{2n})\ngeq 0$, $\Theta$ is neither steering-annihilating nor maximal Gaussian unsteerable.
 So
-${\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$
+${\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}\not\subset{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$
 and
-${\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}).$
+${\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}\not\subset{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}}$.
 There also exist Gaussian steering-annihilating channels which are not Gaussian steering-breaking, that is,
-${\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B}).$
-For example, take $\tilde{\phi}=\phi\otimes I_{B}\in{\mathcal{GC}}(H_{A}\otimes H_{B})$, where $\phi=\phi(\cos\theta I_{2},\sin^{2}\theta I_{2},\mathbf{d}_{0})$ is a single-mode attenuator channel.
+${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}\not\subset{\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}$.
+For example, take $\tilde{\phi}=\phi\otimes I_{B}\in{\mathcal{GC}}^{(1+1)}$, where $\phi=\phi(\cos\theta I_{2},\sin^{2}\theta I_{2},\mathbf{d}_{0})$ is a single-mode attenuator channel.
 Then $\tilde{\phi}=\tilde{\phi}(K,M,\mathbf{d})$ can be represented by
@@
 $=$ $\left(\begin{array}[]{cccc}\sin^{2}\theta&-i\cos^{2}\theta&0&0\\ i\cos^{2}\theta&\sin^{2}\theta&0&0\\ 0&0&0&-i\\ 0&0&i&0\end{array}\right)\ngeq 0.$
+and
+$M+i{\Omega_{2}}=M+i\left(\Omega_{1}\oplus\Omega_{1}\right)$
+$=$ $\left(\begin{array}[]{cccc}\sin^{2}\theta&i&0&0\\ -i&\sin^{2}\theta&0&0\\ 0&0&0&i\\ 0&0&-i&0\end{array}\right)\ngeq 0.$
 Thus, by Theorem 4 and Theorem 7, $\tilde{\phi}$ is steering-annihilating, but not steering-breaking.
@@
 In this section, we discuss two special type of Gaussian unsteerable superchannels, that is, maximal Gaussian unsteerable superchannels and Gaussian unsteerable superchannels.
-Here, we say that a Gaussian superchannel is Gaussian unsteerable if it maps any Gaussian unsteerable channels into Gaussian unsteerable channels; and is maximal Gaussian unsteerable if it maps any maximal Gaussian unsteerable channels into maximal Gaussian unsteerable channels.
-Denote by ${\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ the set of all maximal Gaussian unsteerable superchannels, that is,
-$\begin{array}[]{rl}&{\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\\ =\{&\Phi\in{\mathcal{SGC}}(H_{A}\otimes H_{B}):\\ &\Phi({\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})\},\end{array}$
-and by ${\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$ the set of all Gaussian unsteerable superchannels, that is,
-$\begin{array}[]{rl}&{\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\\ =\{&\Phi\in{\mathcal{SGC}}(H_{A}\otimes H_{B}):\\ &\Phi({\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B}))\subseteq{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})\}.\end{array}$
+Here, we say that a Gaussian superchannel is Gaussian unsteerable from A to B if it maps any Gaussian unsteerable (from A to B) channels into Gaussian unsteerable (from A to B) channels; and is maximal Gaussian unsteerable from A to B if it maps any maximal Gaussian unsteerable (from A to B) channels into maximal Gaussian unsteerable (from A to B) channels.
+Denote by ${\mathcal{GSC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ the set of all $(m+n)$-mode maximal Gaussian unsteerable (from A to B) superchannels, that is,
+$\begin{array}[]{rl}&{\mathcal{GSC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}\\ =\{&\Phi\in{\mathcal{GSC}}^{(m,n)}:\\ &\Phi({\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)})\subseteq{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}\},\end{array}$
+and by ${\mathcal{GSC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ the set of all $(m+n)$-mode Gaussian unsteerable (from A to B) superchannels, that is,
+$\begin{array}[]{rl}&{\mathcal{GSC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}\\ =\{&\Phi\in{\mathcal{GSC}}^{(m,n)}:\\ &\Phi({\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)})\subseteq{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}\},\end{array}$
 To characterize maximal Gaussian unsteerable superchannels, it is necessary to discuss the structure of maximal Gaussian unsteerable channels.
@@
 Theorem 8.
-Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any $(m+n)$-mode Gaussian channel.
-Then $\phi\in{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ if and only if
-$\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(0_{2m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$
+Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{\left(m,n\right)}$ is any $(m+n)$-mode Gaussian channel.
+Then $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ if and only if
+$\mathbf{w}^{{\dagger}}M\mathbf{w}+|\mathbf{w}^{{\dagger}}K\left(0_{2m}\oplus\Omega_{n}\right){K^{\rm T}}\mathbf{w}|$ (18)
 $\geq$ $|\mathbf{w}^{{\dagger}}({0_{2m}}\oplus{\Omega_{n}})\mathbf{w}|$
 holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
-Suppose that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian superchannel.
-By Theorem 1, there exist some ${\chi_{1}}\left(\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},0,0\right),{\chi_{2}}\left(A,Y,\nu\right)\in{\mathcal{GC}}(H_{A}\otimes H_{B})$ such that $\Phi\left(\phi\right)={\chi_{2}}\circ\phi\circ\chi_{1}$ for all $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}(H_{A}\otimes H_{B})$; and moreover, $\Phi\left(\phi\left(K,M,\mathbf{d}\right)\right)=\phi^{\prime}\left(K^{\prime},M^{\prime},\mathbf{d}^{\prime}\right)$ with
-$K^{\prime}=AK\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},\ \ M^{\prime}=AMA^{\rm T}+Y.$
-Take any $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ and any $\rho=\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in{\mathcal{GS}_{\mathcal{US}}(H_{A}\otimes H_{B})}$.
-Notice that the CM $\Gamma_{\tau}$ of any $(m+n)$-mode unsteerable Gaussian state $\tau$ can be written as $\Gamma_{\tau}=0_{2m}\oplus Q_{\tau}+P_{\tau}$ with some real matrix $Q_{\tau}\geq i{\Omega_{n}}$ and $P_{\tau}\geq 0$ MKIB .
-Now, if both $\chi_{1}$ and $\chi_{2}$ are maximal unsteerable, then
-$K^{\prime}\Gamma_{\rho}(K^{\prime})^{\rm T}+M^{\prime}$
-$=$ $A(K\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\Gamma_{\rho}\Sigma_{m+n}E\Sigma_{m+n})K^{\rm T}A^{\rm T}$
-$+AMA^{\rm T}+Y$
-$=$ $A[K(0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)})K^{\rm T}+M]A^{\rm T}+Y$
-$=$ $A(0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)})A^{\rm T}+Y$
-$\geq$ $\pm i(0_{2m}\oplus{\Omega_{n}}),$
-where
-$0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)}=\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\Gamma_{\rho}\Sigma_{m+n}E\Sigma_{m+n},$
-$0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)}=K\left(0_{2m}\oplus\sigma_{\rho}^{(1)}+Q_{\rho}^{(1)}\right)K^{\rm T}+M,$
-the second equation dues to the maximal unsteerability of $\chi_{1}$, the third equation is because $\phi$ is maximal unsteerable, and the last inequality is as $\chi_{2}$ is maximal unsteerable.
-It follows that $\Phi\left(\phi\right)\in{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$.
-On the other hand, suppose that the matrices $A,E$ and $Y$ satisfy
-$Y+\left(0_{2m}\oplus i\Omega_{n}\right)-A\left(0_{2m}\oplus i\Omega_{n}\right){A^{\rm T}}\geq 0$ (18)
-and
-$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}\geq 0.$ (19)
-Noting that $\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}=-\left(0_{2m}\oplus i\Omega_{n}\right)$, Ineq.(19) implies
-$\left(0_{2m}\oplus i\Omega_{n}\right)$ (20)
-$-\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}E\Sigma_{m+n}\geq 0.$
-Comparing Ineqs.(18), (20) and (9) gives $\chi_{1},{\chi_{2}}\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
-In this case, for any $\phi\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$, $K^{\prime}$ and $M^{\prime}$ satisfies
-$M^{\prime}+({0_{2m}}\oplus i{\Omega_{n}})-K^{\prime}\left({0_{2m}}\oplus i{\Omega_{n}}\right){K^{\prime\rm T}}$
-$=$ ${A}MA^{\rm T}+Y+(0_{2m}\oplus i\Omega_{n})$
-$-{A}{K}\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}(0_{2m}\oplus i{\Omega_{n}})\Sigma_{m+n}E\Sigma_{m+n}{{K^{\rm T}}}{{A}^{\rm T}}$
-$\geq$ $AMA^{\rm T}+Y-AK(0_{2m}\oplus i\Omega_{n})K^{\rm T}A^{\rm T}+(0_{2m}\oplus i\Omega_{n})$
-$=$ $A[M+(0_{2m}\oplus i\Omega_{n})-K(0_{2m}\oplus i\Omega_{n})K^{\rm T}]A^{\rm T}$
-$+Y+(0_{2m}\oplus i\Omega_{n})-A(0_{2m}\oplus i\Omega_{n})A^{\rm T}$
-$\geq$ $Y+(0_{2m}\oplus i\Omega_{n})-A(0_{2m}\oplus i\Omega_{n})A^{\rm T}\geq 0.$
-Hence $\Phi\left(\phi\right)={\chi_{2}}\circ\phi\circ{\chi_{1}}\in{\mathcal{GC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
-Based on the above discussions and Theorem 8, we can give sufficient conditions of Gaussian superchannels becoming maximal unsteerable and unsteerable, respectively.
+Thus, by Theorem 8, we can give sufficient conditions of Gaussian superchannels becoming maximal unsteerable and unsteerable, respectively.
 Theorem 9.
-Assume that $\Phi=\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}(H_{A}\otimes H_{B})$ is any $(m+n)$-mode Gaussian superchannel. If either
-(1) there exist maximal Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},0,0)$, $\chi_{2}=\chi_{2}(A,Y,\nu)\in\mathcal{GC}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}(H_{A}\otimes H_{B})$, or
+Assume that $\Phi=\Phi\left(A,E,Y,\nu\right)\in{\mathcal{GSC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian superchannel. If either
+(1) there exist maximal Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(K_{1},M_{1},\mathbf{d}_{1})$, $\chi_{2}=\chi_{2}(K_{2},M_{2},\mathbf{d}_{2})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}^{(m,n)}$, or
 (2) the matrices $A,E$ and $Y$ satisfy
@@
 $\begin{array}[]{rl}&|\mathbf{w}^{{\dagger}}\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\left(0_{2m}\oplus\Omega_{n}\right)\Sigma_{m+n}E\Sigma_{m+n}\mathbf{w}|\\ \geq&|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|\end{array}$
-for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, then $\Phi\in{\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$.
+for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, then $\Phi\in{\mathcal{GSC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$.
 If either
-(3) there exist Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(\Sigma_{m+n}E^{\rm T}\Sigma_{m+n},0,0)$, $\chi_{2}=\chi_{2}(A,Y,\nu)\in\mathcal{GC}_{\mathcal{US}}(H_{A}\otimes H_{B})$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}(H_{A}\otimes H_{B})$, or
-(4) the matrices $A,E$ and $Y$ satisfy Ineqs.(18)-(19), then $\Phi\in{\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
-We remark here that Ineq.(19) is equivalent to the following equation
-$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}=0.$
+(3) there exist Gaussian unsteerabe channels $\chi_{1}=\chi_{1}(K_{1},M_{1},\mathbf{d}_{1})$, $\chi_{2}=\chi_{2}(K_{2},M_{2},\mathbf{d}_{2})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ such that $\Phi(\phi)=\chi_{2}\circ\phi\circ\chi_{1}$ for all $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}^{(m,n)}$, or
+(4) the matrices $A,E$ and $Y$ satisfy
+$Y+\left(0_{2m}\oplus i\Omega_{n}\right)-A\left(0_{2m}\oplus i\Omega_{n}\right){A^{\rm T}}\geq 0$
+and
+$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}\geq 0,$
+then $\Phi\in{\mathcal{GSC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.
+Proofs of Theorems 8 and 9 are given in Appendix.
 Notice that the converse of Theorem 9 may not be true.
@@
 In addition, there exists maximal Gaussian unsteerable superchannels which are not Gaussian unsteerable superchannels.
-For example, take a $(2+2)$-mode Gaussian superchannel $\Phi=\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}(H_{A}\otimes H_{B})$ with
+For example, take a $(1+1)$-mode Gaussian superchannel $\Phi=\Phi\left(A,E,Y,\nu\right)$ with
 […]
@@
 By calculations, we find that $A,E,Y$ satisfy
-${w^{\dagger}}Yw+\left|{{w^{\dagger}}A\left({{0_{A}}\oplus{\Omega_{B}}}\right){A^{T}}w}\right|\geq\left|{{w^{\dagger}}\left({{0_{A}}\oplus{\Omega_{B}}}\right)w}\right|$}
+${w^{\dagger}}Yw+\left|{{w^{\dagger}}A\left({{0_{2}}\oplus{\Omega_{1}}}\right){A^{\rm T}}w}\right|\geq\left|{{w^{\dagger}}\left({{0_{2}}\oplus{\Omega_{1}}}\right)w}\right|$}
 for any $w\in\mathbb{C}^{4}$, but do not satisfy
-$Y-i\left({{0_{A}}\oplus{\Omega_{B}}}\right)+A\left({{0_{A}}\oplus i{\Omega_{B}}}\right){A^{T}}\geq 0.$}
-That implies that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B})$, but $\Phi\left(A,E,Y,\nu\right)\notin{\mathcal{SGC}}_{\mathcal{US}}(H_{A}\otimes H_{B})$.
+$Y-i\left({{0_{2}}\oplus{\Omega_{1}}}\right)+A\left({{0_{2}}\oplus i{\Omega_{1}}}\right){A^{\rm T}}\geq 0.$}
+That implies that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{GSC}}^{(1,1)}_{\mathcal{MUS}(A\rightarrow B)}$, but $\Phi\left(A,E,Y,\nu\right)\notin{\mathcal{GSC}}^{(1,1)}_{\mathcal{US}(A\rightarrow B)}$.
+Thus, the set of Gaussian maximal unsteerable superchannels differs from that of all Gaussian unsteerable superchannels.
+Nevertheless, both can be identified as free superchannels within the resource theory of steering for channels in CV systems.
 V Conclusion
@@
 VI Acknowledments
-This work is partially supported by National Natural Science Foundation of China (11901421, 12571138, 12171290) and the special fund for Science and Technology Innovation Teams of Shanxi Province (202304051001035).
+The authors wish to give their thanks to the referees for many helpful comments to improve the original paper.
+This work is partially supported by National Natural Science Foundation of China (11901421, 12571138, 12171290), Shanxi Scholarship Council of China (2025-001) and the special fund for Science and Technology Innovation Teams of Shanxi Province (202304051001035).
+Appendix
+A proof of Theorem 6.
+Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$ is any $(m+n)$-mode Gaussian steering-annihilating channel.
+Then for any $\rho\in{\mathcal{GS}}^{(m,n)}$ with CM $\Gamma_{\rho}$, $\phi(\rho)$ is unsteerable from A to B or from B to A.
+If $\phi(\rho)\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$, by Eqs.(5)-(6), one has
+$K\Gamma_{\rho}K^{\rm T}+M\geq\pm i(0_{2m}\oplus\Omega_{n}),$
+and so
+$\mathbf{w}^{{\dagger}}(K\Gamma_{\rho}K^{\rm T}+M)\mathbf{w}\geq\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w},$
+that is,
+$\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}$
+holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
+Note that, it is shown in PMGH that, for any $N$-mode Gaussian state $\rho$ with CM $\Gamma$, it holds that
+$\inf\limits_{\Gamma\geq\pm i\Omega_{N}}{\bf w}^{\dagger}\Gamma{\bf w}=|{\bf w}^{\dagger}\Omega_{N}{\bf w}|,\ \ \forall\mathbf{w}\in\mathbb{C}^{2N}.$ (19)
+Thus, by Eq.(19), we have
+$\begin{array}[]{rl}&\inf\limits_{\Gamma_{\rho}\geq\pm i(\Omega_{m}\oplus\Omega_{n})}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ =&|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ \geq&\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w},\end{array}$
+which implies
+$|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
+for all $\mathbf{w}\in\mathbb{C}^{2(m+n)},$ that is, Ineq.(12) holds.
+Symmetrically, one can show that, if $\phi(\rho)\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(B\rightarrow A)}$, then Ineq.(13) holds.
+On the other hand, if Ineq.(12) holds, then for any $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, by using Eq.(19), one has
+$\mathbf{w}^{{\dagger}}\Gamma_{\phi(\rho)}\mathbf{w}$ $=$ $\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
+$\geq$ $\inf_{\Gamma_{\rho}\geq\pm i\left({\Omega_{m}\oplus{\Omega_{n}}}\right)}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
+$=$ $|\mathbf{w}^{{\dagger}}K(\Omega_{m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}$
+$\geq$ $|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
+$\geq$ $\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.$
+So $\Gamma_{\phi(\rho)}\geq\pm i(0_{2m}\oplus\Omega_{n})$, which means that $\phi(\rho)$ is unsteerable from A to B.
+Symmetrically, if Ineq.(13) holds, then $\phi(\rho)$ is unsteerable from B to A.
+It follows that $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$.
+∎
+A proof of Theorem 7.
+Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{\left(N\right)}$ is any $N$-mode Gaussian channel.
+$\left(1\right)\Rightarrow\left(2\right)$: By Definition 2, this is obvious.
+$\left(2\right)\Rightarrow\left(3\right)$: Note that the CM $\Gamma_{{\rho_{\phi}}}$ of $\rho_{\phi}$} has the form X11
+[…]
+If ${\rho_{\phi}}\in{\mathcal{GS}}^{(N,N)}_{\mathcal{US}(A\rightarrow B)}$, then
+$\Gamma_{{\rho_{\phi}}}+(0_{2N}\oplus i{\Omega_{N}})\geq 0.$ (21)
+It is well known ZH that a Hermitian matrix
+$W=\left(\begin{array}[]{cc}W_{11}&W_{12}\\ W_{12}^{{\dagger}}&W_{22}\end{array}\right)\geq 0$ (24)
+$\Leftrightarrow W_{22}\geq 0,\ \ W_{11}-W_{12}W_{22}^{-1}W_{12}^{{\dagger}}\geq 0$
+$\Leftrightarrow W_{11}\geq 0,\ \ W_{22}-W_{12}^{{\dagger}}W_{11}^{-1}W_{12}\geq 0.$
+So Ineq.(21) implies
+$\begin{cases}\cosh 2rI_{2N}+i{\Omega_{N}}\geq 0,\\ \begin{array}[]{rl}&\cosh 2rK{K^{\rm T}}+M\\ &-\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N}+i\Omega_{N})^{-1}\Sigma_{N}{K^{\rm T}}\geq 0.\end{array}\end{cases}$
+Note that $(T+cI)^{-1}=\frac{1}{c}(I-\frac{T}{c})+o(c^{-2})$ S1 , where $o(c^{-2})$ stands for the Landau little-$o$ which will be neglected when taking to $c\to\infty$. So
+$0\leq\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N}+i{\Omega_{N}})^{-1}\Sigma_{N}K^{\rm T}]$
+$=$ $\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-\sinh^{2}(2r)K\Sigma_{N}[\frac{1}{\cosh 2r}(I_{2N}-\frac{i\Omega_{N}}{\cosh 2r})+o(\frac{1}{\cosh 2r})]\Sigma_{N}{K^{\rm T}}]$
+$=$ $\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M-K\Sigma_{N}[\frac{\sinh^{2}(2r)}{\cosh 2r}I_{2N}-i\tanh^{2}(2r){\Omega_{N}}+\sinh^{2}(2r)o(\frac{1}{\cosh 2r})]\Sigma_{N}K^{\rm T}]=M-iK{\Omega_{N}}{K^{\rm T}}.$
+If ${\rho_{\phi}}\in{\mathcal{GS}}^{(N,N)}_{\mathcal{US}(B\rightarrow A)}$, then
+$\Gamma_{{\rho_{\phi}}}+(i\Omega_{N}\oplus 0_{2N})\geq 0.$ (25)
+By the equivalence (24) again, Ineq.(25) implies
+$\begin{array}[]{rl}&\cosh 2rK{K^{\rm T}}+M+i\Omega_{N}\\ -&\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N})^{-1}\Sigma_{N}{K^{\rm T}}\geq 0.\end{array}$
+So
+$0\leq\mathop{\lim}\limits_{r\to\infty}[\cosh 2rK{K^{\rm T}}+M+i\Omega_{N}$
+- $\sinh^{2}(2r)K\Sigma_{N}(\cosh 2rI_{2N})^{-1}\Sigma_{N}{K^{\rm T}}]$
+$=$ $\mathop{\lim}\limits_{r\to\infty}[\frac{\cosh^{2}(2r)-\sinh^{2}(2r)}{\cosh 2r}K{K^{\rm T}}+M+i\Omega_{N}]$
+$=$ $M+i{\Omega_{N}}.$
+Therefore, the statement (2) holds.
+[…]
+[…]
+Since $\Gamma_{\rho}+i({\Omega_{N}}\oplus{\Omega_{N^{\prime}}})\geq 0$, a direct calculation gives
+$Y+i{\Omega_{N^{\prime}}}\geq 0$}
+and
+$X+i{\Omega_{N}}-Z\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}{Z^{\rm T}}\geq 0$}.
+If the matrices $K$ and $M$ satisfy $M-iK\Omega_{N}K^{\rm T}\geq 0$, one gets
+$KX{K^{\rm T}}+M-KZ{\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}}{Z^{\rm T}}{K^{\rm T}}$
+$=$ $K\left({X-Z{{\left({Y+i{\Omega_{N^{\prime}}}}\right)}^{-1}}{Z^{\rm T}}}\right){K^{\rm T}}+M$
+$=$ $K\left({X-Z{{\left({Y+i{\Omega_{N^{\prime}}}}\right)}^{-1}}{Z^{\rm T}}}+i{\Omega_{N}}\right){K^{\rm T}}$
+$+(M-iK{\Omega_{N}}{K^{\rm T}})\geq 0.$
+Also note that $KX{K^{\rm T}}+M\geq 0$ by the denifition of the channel $\phi$. By ZH again, one achieves
+${\Gamma_{\left({\phi\otimes I_{H^{\prime}}}\right)(\rho)}}+i(0_{N}\oplus{\Omega_{N^{\prime}}})\geq 0.$
+So $(\phi\otimes{I_{H^{\prime}}})\left(\rho\right)$ is unsteerable from A to B, and hence $\phi$ is steering-breaking.
+Now, assume that the matrices $K$ and $M$ satisfy the condition $M+i\Omega_{N}\geq 0$.
+It is shown that in LHQ that, if the matrices $L,V,W\in\mathcal{M}_{n}(\mathbb{C})$ satisfy the condition $L>0$, $W\geq 0$ and $VLV^{{\dagger}}+W$ invertible, then
+$V^{{\dagger}}(VLV^{{\dagger}}+W)^{-1}V\leq L^{-1}.$
+For $X$, $K$ and $M$, applying the above result, one achieves
+${K^{\rm T}}(KX{K^{\rm T}}+(M+i{\Omega_{N}}))^{-1}K\leq X^{-1}.$
+and so
+$Y-{Z^{\rm T}}{K^{\rm T}}(KX{K^{\rm T}}+M+i{\Omega_{N}})^{-1}KZ$
+$\geq$ $Y-{Z^{\rm T}}X^{-1}{Z}\geq 0.$
+By the equivalence (24), one obtains
+${\Gamma_{\left({\phi\otimes I_{H^{\prime}}}\right)(\rho)}}+i({\Omega_{N}}\oplus 0_{N}^{\prime})\geq 0.$
+So $(\phi\otimes{I_{H^{\prime}}})\left(\rho\right)$ is unsteerable from B to A. Hence $\phi$ is steering-breaking. ∎
+A proof of Theorem 8.
+Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{\left(m,n\right)}$ is any $(m+n)$-mode Gaussian channel and $\rho\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ is any $(m+n)$-mode Gaussian unsteerable state with CM $\Gamma_{\rho}$.
+On the one hand, if $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$, then $\phi(\rho)\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$. So
+$K\Gamma_{\rho}K^{\rm T}+M\geq\pm i(0_{2m}\oplus\Omega_{n}),$
+which implies
+$\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}$
+for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$. By a similar discussion to that of Eq.(19), one can obtain
+$\begin{array}[]{rl}&\inf\limits_{\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n})}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}K^{\rm T}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ =&|{\bf w}^{\dagger}K(0_{2m}\oplus\Omega_{n})K^{\rm T}{\bf w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\\ \geq&\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.\end{array}$
+Hence
+$|\mathbf{w}^{{\dagger}}K(0_{2m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}\geq|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
+holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
+On the other hand, if (18) holds, as $\Gamma_{\rho}\geq\pm i\left({0_{2m}\oplus{\Omega_{n}}}\right)$, for any $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, we have
+$\mathbf{w}^{{\dagger}}\Gamma_{\phi(\rho)}\mathbf{w}$ $=$ $\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
+$\geq$ $\inf_{\Gamma_{\rho}\geq\pm i(0_{2m}\oplus\Omega_{n})}\mathbf{w}^{{\dagger}}K\Gamma_{\rho}{K^{\rm T}}\mathbf{w}+\mathbf{w}^{{\dagger}}M\mathbf{w}$
+$=$ $|\mathbf{w}^{{\dagger}}K(0_{2m}\oplus\Omega_{n})K^{\rm T}\mathbf{w}|+\mathbf{w}^{{\dagger}}M\mathbf{w}$
+$\geq$ $|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
+$\geq$ $\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.$
+This means $\Gamma_{\phi(\rho)}\geq\pm i(0_{2m}\oplus\Omega_{n})$, and so $\phi(\rho)$ is unsteerable from A to B. It follows that $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$. ∎
+A proof of Theorem 9.
+Suppose that $\Phi\left(A,E,Y,\nu\right)\in{\mathcal{SGC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian superchannel.
+By X11 , there exist some $\chi_{1}=\chi_{1}(K_{1},M_{1},\mathbf{d}_{1})$, $\chi_{2}=\chi_{2}(K_{2},M_{2},\mathbf{d}_{2})\in{\mathcal{GC}}^{(m,n)}$ such that $\Phi\left(\phi\right)={\chi_{2}}\circ\phi\circ\chi_{1}=\phi^{\prime}(K^{\prime},M^{\prime},\mathbf{d})$ for all $\phi\left(K,M,\mathbf{d}\right)\in\mathcal{GC}^{(m,n)}$.
+Note that $K^{\prime}$ and $M^{\prime}$ satisfy
+$M^{\prime}={K_{2}}K{M_{1}}{K^{\rm T}}{K_{2}}^{\rm T}+{K_{2}}M{K_{2}}^{\rm T}+{M_{2}}$
+and
+$K^{\prime}=K_{2}{K}K_{1}.$
+Also notice that the CM $\Gamma_{\tau}$ of any $(m+n)$-mode unsteerable (from A to B) Gaussian state $\tau$ can be written as
+$\Gamma_{\tau}=0_{2m}\oplus Q_{\tau}+P_{\tau}$
+with some real matrix $Q_{\tau}\geq i{\Omega_{n}}$ and $P_{\tau}\geq 0$ MKIB .
+Now, if both $\chi_{1}$ and $\chi_{2}$ are maximal unsteerable from A to B, then for any $(m+n)$-mode maximal Gaussian unsteerable channel $\phi=\phi\left(K,M,\mathbf{d}\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$ and any Gaussian unsteerable state $\rho=\rho(\Gamma_{\rho},\mathbf{d}_{\rho})\in{\mathcal{GS}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$, we have
+$\Gamma_{\Phi\left(\phi\right)(\rho)}=\Gamma_{{\chi_{2}}\circ\phi\circ{\chi_{1}}(\rho)}$
+$=$ ${K}_{2}K{K_{1}}\Gamma_{\rho}{K_{1}}^{\rm T}{K^{\rm T}}{K_{2}}^{\rm T}$
+$+{K_{2}}K{M_{1}}{K^{\rm T}}{K_{2}}^{\rm T}+{K_{2}}M{K_{2}}^{\rm T}+{M_{2}}$
+$=$ ${K}_{2}[K(0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)})K^{\rm T}+M]{K}_{2}^{\rm T}+M_{2}$
+$=$ ${K}_{2}(0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)}){K}_{2}^{\rm T}+M_{2}$
+$\geq$ $\pm i(0_{2m}\oplus{\Omega_{n}}),$
+where
+$0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)}=K_{1}\Gamma_{\rho}K_{1}^{\rm T}+M_{1},$
+$0_{2m}\oplus Q_{\rho}^{(2)}+P_{\rho}^{(2)}=K\left(0_{2m}\oplus Q_{\rho}^{(1)}+P_{\rho}^{(1)}\right)K^{\rm T}+M,$
+and the second equation, the third equation, the last inequality due to respectively the maximal unsteerability of $\chi_{1}$, $\phi$ and $\chi_{2}$.
+So $\Phi(\phi)(\rho)$ is unsteerable from A to B.
+It follows that $\Phi\left(\phi\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}$.
+If both $\chi_{1}$ and $\chi_{2}$ are Gaussian unsteerable from A to B, that is, they satisfy Ineq.(9):
+$M_{2}+\left(0_{2m}\oplus i\Omega_{n}\right)-K_{2}\left(0_{2m}\oplus i\Omega_{n}\right){K_{2}^{\rm T}}\geq 0$
+and
+$M_{1}+\left(0_{2m}\oplus i\Omega_{n}\right)-K_{1}\left(0_{2m}\oplus i\Omega_{n}\right){K_{1}^{\rm T}}\geq 0,$
+then for any $(m+n)$-mode Gaussian unsteerable channel $\phi({K,M,\mathbf{d}})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$, we have
+$M+\left(0_{2m}\oplus i\Omega_{n}\right)-K\left(0_{2m}\oplus i\Omega_{n}\right){K^{\rm T}}\geq 0,$
+and so
+$M^{\prime}+({0_{2m}}\oplus i{\Omega_{n}})-K^{\prime}\left({0_{2m}}\oplus i{\Omega_{n}}\right){K^{\prime\rm T}}$
+$=$ ${K_{2}}K{M_{1}}{K^{\rm T}}{K_{2}}^{\rm T}+{K_{2}}M{K_{2}}^{\rm T}+{M_{2}}$
+$+({0_{2m}}\oplus i{\Omega_{n}})-K_{2}{K}K_{1}(0_{2m}\oplus i{\Omega_{n}})K_{1}^{\rm T}K^{\rm T}{{K_{2}}^{\rm T}}$
+$=$ ${K_{2}}K[M_{1}+({0_{2m}}\oplus i{\Omega_{n}})-K_{1}(0_{2m}\oplus i{\Omega_{n}})K_{1}^{\rm T}]K^{\rm T}{K_{2}}^{\rm T}$
+$+{K_{2}}[M+({0_{2m}}\oplus i{\Omega_{n}})-K(0_{2m}\oplus i{\Omega_{n}})K^{\rm T}]{K_{2}}^{\rm T}$
+$+{M_{2}}+({0_{2m}}\oplus i{\Omega_{n}})-K_{2}(0_{2m}\oplus i{\Omega_{n}}){{K_{2}}^{\rm T}}$
+$\geq$ $0.$
+Hence $\Phi\left(\phi\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.
+On the other hand, suppose that the matrices $A,E$ and $Y$ satisfy
+$Y+\left(0_{2m}\oplus i\Omega_{n}\right)-A\left(0_{2m}\oplus i\Omega_{n}\right){A^{\rm T}}\geq 0$
+and
+$\left(0_{2m}\oplus i\Omega_{n}\right)-E\left(0_{2m}\oplus i\Omega_{n}\right)E^{\rm T}\geq 0.$
+As $\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}=-\left(0_{2m}\oplus i\Omega_{n}\right)$, the inequality implies
+$\begin{array}[]{rl}&\left(0_{2m}\oplus i\Omega_{n}\right)\\ &-\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}\left(0_{2m}\oplus i\Omega_{n}\right)\Sigma_{m+n}E\Sigma_{m+n}\geq 0.\end{array}$
+Thus, for any $\phi(K,M,\mathbf{d})\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ with $\Phi(\phi)=\phi^{\prime}(K^{\prime},M^{\prime},\mathbf{d}^{\prime})$, $K^{\prime}$ and $M^{\prime}$ satisfy
+$M^{\prime}+({0_{2m}}\oplus i{\Omega_{n}})-K^{\prime}\left({0_{2m}}\oplus i{\Omega_{n}}\right){K^{\prime\rm T}}$
+$=$ ${A}MA^{\rm T}+Y+(0_{2m}\oplus i\Omega_{n})$
+$-{A}{K}\Sigma_{m+n}E^{\rm T}\Sigma_{m+n}(0_{2m}\oplus i{\Omega_{n}})\Sigma_{m+n}E\Sigma_{m+n}{{K^{\rm T}}}{{A}^{\rm T}}$
+$\geq$ $AMA^{\rm T}+Y-AK(0_{2m}\oplus i\Omega_{n})K^{\rm T}A^{\rm T}+(0_{2m}\oplus i\Omega_{n})$
+$=$ $A[M+(0_{2m}\oplus i\Omega_{n})-K(0_{2m}\oplus i\Omega_{n})K^{\rm T}]A^{\rm T}$
+$+Y+(0_{2m}\oplus i\Omega_{n})-A(0_{2m}\oplus i\Omega_{n})A^{\rm T}$
+$\geq$ $Y+(0_{2m}\oplus i\Omega_{n})-A(0_{2m}\oplus i\Omega_{n})A^{\rm T}\geq 0.$
+Hence $\Phi\left(\phi\right)\in{\mathcal{GC}}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$ in this case. ∎